### v1

# https://github.com/yfzhang114/LLaVA-Align Debiasing Large Visual Language Models

Abstract
In the realms of computer vision and natural language processing, Large Vision-Language Models (LVLMs) have become indispensable tools, proficient in generating textual descriptions based on visual inputs. Despite their advancements, our investigation reveals a noteworthy bias in the generated content, where the output is primarily influenced by the underlying Large Language Models (LLMs) prior rather than the input image. Our empirical experiments underscore the persistence of this bias, as LVLMs often provide confident answers even in the absence of relevant images or given incongruent visual input. To rectify these biases and redirect the model’s focus toward vision information, we introduce two simple, training-free strategies. Firstly, for tasks such as classification or multi-choice question-answering (QA), we propose a “calibration” step through affine transformation to adjust the output distribution. This “Post-Hoc debias” approach ensures uniform scores for each answer when the image is absent, serving as an effective regularization technique to alleviate the influence of LLM priors. For more intricate open-ended generation tasks, we extend this method to “Debias sampling”, drawing inspirations from contrastive decoding methods. Furthermore, our investigation sheds light on the instability of LVLMs across various decoding configurations. Through systematic exploration of different settings, we significantly enhance performance, surpassing reported results and raising concerns about the fairness of existing evaluations. Comprehensive experiments substantiate the effectiveness of our proposed strategies in mitigating biases. These strategies not only prove beneficial in minimizing hallucinations but also contribute to the generation of more helpful and precise illustrations.
Keywords:
Large Vision-Language Models Debiasing Hallucination1 Introduction
LLMs such as GPT-4 [34], PaLM [7], and LLaMA-2 [47] have evolved to handle both text and images. This expansion into the multimodal domain involves further pre-training with image-text pairs or fine-tuning with specialized vision instruction tuning datasets, resulting in the birth of powerful Large Visual Language Models (LVLMs). These models excel at understanding complex visual patterns and translating them into coherent language. However, despite their capabilities, our investigation reveals a notable issue. The content generated by LVLMs is significantly biased towards the underlying LLMs used during pre-training, rather than being influenced by the input images. Even in situations where the image is entirely noisy or absent, LVLMs confidently generate answers, indicating a bias towards the learned language patterns. This observation is illustrated in Fig. 1, where querying an LVLM about the color of a non-existent fox produces confident yet unrealistic answers like brown, gray, or blue. To explore this bias further, we conduct a toy experiment using questions about color, shape, number, and relationships for 80 entities in the MSCOCO dataset. Our experimental baselines include Noise (fully noisy image and question input), None (no image provided), and Zeros/Ones (image with all zero/one values). Fig. 1 displays the top 15 choices and corresponding probabilities of frequently occurring answers across these cases. Remarkably, even when no image is available or questions are about non-existent or meaningless images, current LVLMs tend to produce specific answers. This indicates a substantial bias originating from the pretraining on LLMs. Such biases bring persistent challenges, particularly the issue of hallucination [13, 26, 26, 43], pose a significant concern, impacting the reliability and applicability of LVLMs. In this paper, we also aim to uncover the underlying reasons behind such a phenomenon, and the results will provide insights into the design of current LVLMs and their training strategies.
In this paper, we present two training-free strategies to mitigate generation biases associated with underlying LLMs. Our first approach introduces a “calibration” step to adjust the output distribution. Specifically, we assess the model’s inclination toward specific answers by inputting a dummy test with no image, as illustrated in Fig. 1, where we replace the image with ’None’ or meaningless tensors. We then determine the calibration parameters to ensure that the image-free input produces uniform scores for each answer without yielding overly confident outputs. This calibration procedure establishes suitable parameters without requiring additional training data, referred to as the ‘‘Post-Hoc debias’’ approach, proving intuitive for classification tasks111In binary classification settings like POPE, we aim for the input with a meaningless image to have ’yes’ and ’no’ with an equal probability of .. However, in open-ended generation, calibrating each token’s logits uniformly, even without image input, is unreasonable due to the inherent correlation between text tokens. In such cases, we extend the Post-Hoc debias method to its ‘‘Debias sampling’’ counterpart. Inspired by contrastive decoding methods [25, 21], we compute the difference between generation token log-probabilities for one with a correct image and another with a meaningless vision input. Both the Post-Hoc debias and debias sampling methods serve as regularization techniques, reducing the generation results’ dependency on pure text or meaningless image input.
Moreover, we have identified a significant level of instability in the performance of current LVLMs across various generative configurations. Our primary hypothesis is that existing evaluations are predominantly based on a default decoding setting, limiting the exploration of the model’s full capabilities. As extensively illustrated in Fig. 2, diverse choices in generative configurations yield substantially different performance outcomes. Notably, adjusting the generation temperature can lead to a notable increase in the average F1 score of POPE-MSCOCO, shifting from 76 to approximately 84.04. This improvement surpasses the previous results achieved by the 13B model and the RLHF-tuned model by a considerable margin, indicating that the current estimation of LVLM capability may be underestimated. Furthermore, our investigation reveals that different models exhibit preferences for distinct generation configurations, highlighting the instability in the current evaluation of LVLMs. This raises concerns about the fairness of evaluations, particularly since they often rely on default generation configurations or selectively choose the best configuration for the proposed model. To maximize the full potential of existing LVLMs, we systematically investigate six LVLM models across four benchmarks, meticulously searching their generative configurations to identify optimal settings. Our results significantly outperform previously reported ones by a substantial margin, underscoring the importance of selecting the best decoding configuration for each model.
2 Method
Background of Large Vision-Language Models. A LVLM with parameters , conditioned on a query text , a vision input , parametrizes a probability distribution over response sequences . Each response token is sampled from a token probability distribution, namely , where denotes the logits of . Usually, an LLM is first pre-trained on a large, unlabeled text dataset, then an LVLM utilizes the LLM as the text encoder and incorporates vision representations in different ways [3, 23, 9, 12]. Related work about LVLMs and Hallucination in LVLMs can be found in Appendix 0.A.
2.1 Exploring the Impact of LLM Biases on LVLMs
Addressing Biases in Current LVLMs. Despite the acknowledged influence of language priors on VQA tasks [15, 57, 53], previous efforts have primarily concentrated on mitigating biases within the training sets. These approaches involve designing advanced training strategies and constructing balanced datasets. However, extending such methods to LVLMs, which utilize pre-existing LLMs as text encoders, is challenging due to the complexity and cost associated with retraining or fine-tuning. Furthermore, it remains unclear whether LVLM performance is affected by the language priors ingrained in their used LLMs.
To assess the influence of LLM biases on LVLMs, we conduct toy experiments with five target questions detailed in Tab. 4. For each entity of the MSCOCO dataset (appendix Tab. 3), we design corresponding prompts and pair them with different types of vision inputs, including unrelated vision input such as fully noisy images (Noise) or entirely black or white images (Zero/One), or replace vision input with pure text, where we either remove all vision tokens (None) or replace them with a meaningless placeholder like </unk> (Unk). As illustrated in Fig. 1 and in the appendix Fig. 13, LVLMs exhibit biases towards specific answers, such as associating “white” with color-related questions, “round” with shape-related questions, and “many” with number-based questions. We further analyze attention mechanisms to explain why models heavily rely on LLM biases. Fig. 4 and 14 demonstrate that LLM outputs tend to allocate more attention to text tokens. Even for original image-text pairs where the number of image tokens exceeds that of text tokens, the sum of attention scores for text surpasses . An intriguing observation is that when we feed the original text-image pair (Naive), pure text (None), and fully noisy images (Noise), the model exhibits similar attention patterns. Most attention is directed towards special but uninformative tokens, rather than vision tokens or text sequences, such as the question itself. In open-ended generation tasks, we randomly select two questions from LLaVA-Bench (as shown in Fig. 11) and examine the attention scores as more tokens were generated. As depicted in Fig. 5, with the increasing length of generated text during open-ended tasks, there was a corresponding reduction in attention allocated to images, exacerbating the issue. Consequently, the model becomes increasingly prone to generating content independently of input images, shedding light on the genesis of hallucinated content from an attention perspective. Additionally, the results also indicate that the shallow layers assign higher attention to visual tokens; however, as the feature progresses into deeper layers, attention on vision tokens will be reduced.
Underlying Reasons: Several factors contribute to the existence of this phenomenon, with the modality gap [27] standing out as a pivotal influencer. This gap becomes apparent as image-text embeddings consistently reside in distinct spaces with substantial distances, a characteristic observed even in models trained on a significant number of image-text pairs [37]. Unfortunately, LVLMs consistently undergo training with a restricted vision-language dataset [43], resulting in a weakened feature alignment. In contrast, the LLMs involved in the process benefit from a more extensive training corpus and demonstrate heightened proficiency in addressing text-based tasks. The pronounced modality gap guides LLMs to predominantly focus on familiar tokens or representations, often neglecting crucial vision tokens. Additionally, following the shallow-to-deep principle[39, 35], where shallower networks can adeptly adapt to changes in data streams or learn more efficiently with limited data, we can explain why shallow layers exhibit better attention score assignments compared to deeper layers. Shallow layers may be better tuned, but deeper layers retain most of the original LLM patterns and tend to ignore unfamiliar vision tokens. Furthermore, LVLMs may inherit the drawbacks of LLMs, which often overlook contextual information and heavily rely on their prior parametric knowledge[42, 5, 54], further diminishing the impact of vision tokens. These revelations underscore inherent limitations in current training strategies and datasets, emphasizing the need for enhanced alignment between the two modalities or a heightened recognition of the contribution made by vision tokens. In response to this challenge, we present a series of training-free methods in this paper, with the aim of mitigating the observed phenomenon and elevating the overall performance of LVLMs.
2.2 Model Calibration
Post-hoc Debiasing of Prediction Results (Fig. 3 (left)). In addressing the bias issue, we implement a “calibration” step for the model’s output probabilities using an affine transformation [36, 14, 52]: . In this equation, a weight matrix and a bias vector are applied to the original probabilities, producing adjusted probabilities. For classification tasks, represents the probabilities associated with each label name, normalized to sum up to one. To enhance efficiency, we constrain the matrix to be diagonal, utilizing vector scaling [14, 52]. An intuitive approach to learning is to initially estimate the LLM’s bias towards certain answers by inputting a meaningless vision input , such as None or Noise. The resulting output prediction is denoted as . Ideally, LLMs should assign a uniform distribution score to this test input. For instance, when presented with the question “Is the color of the fox in the image brown?”, supplying a meaningless image or no image should result in equal probabilities for ’yes’ and ’no’. This is because the model cannot make a decision. However, due to the model’s biases, it tends to assign a higher score to ’yes’. This error can be corrected by setting and to an all-zero vector. Subsequently, the debiased result is obtained as . In our experimental findings, we establish that employing unrelated vision inputs (Noise, Zero, One) does not consistently enhance model performance and exhibits sensitivity to question types. Consequently, as a default debiasing approach, we opt for image-free methods, specifically None and Unk. Additionally, we explore a debiasing method using Both image-free inputs. In this configuration, we set to None and , obtaining respective probability distributions and . We then calculate the probability distribution .
Visual Debias Decoding (VDD) (Fig. 3 (Right)). While post-hoc debias methods can be naturally extended to open-ended generation settings with the entire vocabulary as the label space, implementation faces challenges due to the large label space. At the same time, it’s crucial to maintain the correlation or co-occurrence of output text tokens and simply enforcing a uniform LM output distribution could significantly harm generation quality. In response, we adopt the concept of contrastive decoding [25, 21, 50, 42] and introduce the VDD strategy. Similar to post-hoc debiasing, we feed the original image and text into the LVLM to obtain logits and feed image-free input to get logits , where contains only text priors. To emphasize the contribution of visual information, we aim to mitigate undesired behaviors highlighted by the image-free logits and generate text based on the remaining positive behaviors of the LVLM when provided with the image input. To implement this concept, we propose the contrastive objective: . In this context, represents the amplification level of the debiasing method. It is crucial to recognize that the logits from pure-text models are not invariably incorrect. They can adeptly capture various fundamental aspects of English grammar and common sense. Therefore, adopting a universal penalty for all tokens may not be suitable in every instance. To tackle this nuanced challenge, we introduce an adaptive plausibility constraint across the output vocabulary of LVLMs. This constraint is intricately linked to the confidence level associated with the output distribution when original visual inputs are considered: Here, is a hyperparameter that truncates the next token distribution of . A larger entails more aggressive truncation, keeping only high probability tokens. During generation, we set the logits of all tokens not in to :
The adaptive plausibility constraint adeptly reserves high-probability tokens considering vision information, mitigating bias through the utilization of pure-text logits. This constraint maintains as few as one token in the candidate pool when the expert exhibits high confidence in a specific token, effectively mitigating the influence of the contrastive objective. Importantly, our VDD approach only necessitates the input of pure text into LVLMs, enhancing efficiency compared to visual contrastive decoding [32]. This avoids the intricacies of VCD associated with employing a distorted image to generate reference logits, a process that requires meticulous adjustment of the injected noise level into the image. Similar to post-hoc debiasing methods, here we denote the methods as VDD-None, VDD-Unk, and VDD-Both, corresponding to the use of image-free inputs None, Unk, and both for obtaining reference logits.
2.3 Impact of Decoding Configuration on Model Performance
We observe that existing open-source LVLMs undergo evaluations using a variety of generation strategies, as outlined in Tab. 5. This variability in evaluation configuration introduces a potential source of performance variability, impacting the thorough exploration of LVLM capabilities. Our research primarily delves into decoding configurations, specifically experimenting with distinct strategies:
1. Temperature Sampling: We systematically vary the temperature parameter () from 0.05 to 1, with a step size of 0.05, resulting in 20 configurations. The temperature parameter modulates the sharpness of the next-token distribution.
2. Top- Sampling: By filtering the K most likely next words, where K varies within 1, 2, 5, 10, 20, 50, 100, 200, 500, we generate 9 configurations.
3. Top- Sampling: We select words from the smallest set whose cumulative probability exceeds the threshold . The threshold ranges from 0.05 to 1, with a step size of 0.05, yielding 20 configurations.
For each test sample of a given benchmark, the model generates 49 responses, corresponding to each decoding configuration. In the experimental section, we establish four settings: Temp , Top-, and Top-, where the best response for each test sample is selected among configurations within these three settings. Finally, the Overall setting chooses the best response from all 49 generated responses for each test sample. This systematic exploration of decoding strategies contributes to a more comprehensive understanding of LVLM performance.
3 Experiments
We conduct a series of extensive experiments to demonstrate that: (i) the proposed debiasing methods significantly alleviate hallucination and improve reasoning capability; (ii) exploring the optimal generation configuration unleashes the full potential of existing LVLMs, resulting in substantial performance improvements compared to default configurations; (iii) the debiasing methods rectify model predictions, particularly in situations where the model lacks confidence and is prone to errors; and (iv) we scrutinize the failure cases of our methods, revealing certain flaws in current evaluation benchmarks.
3.1 Datasets and baselines
The paper employs a comprehensive evaluation strategy, leveraging four distinct datasets to assess the proficiency of LVLMs. Firstly, the MMMU dataset, spanning six core disciplines, rigorously tests models on multimodal questions extracted from college exams, quizzes, and textbooks, emphasizing advanced perception and reasoning [51]. Secondly, the MME dataset systematically evaluates perception and cognition capabilities, encompassing tasks like OCR, object recognition, and reasoning across diverse modalities and cognitive domains [11]. The subsets within MME, such as MME-OH and MME-AH, specifically target object-level and attribute-level hallucination assessments [21]. Additionally, MME-RE includes reasoning subsets like code reasoning, text translation, commonsense reasoning, and numerical calculation to gauge cognitive abilities. Thirdly, the POPE benchmark introduces a refined approach for evaluating object hallucination in LVLMs. This benchmark underscores balanced sampling settings and draws data from MSCOCO, A-OKVQA, and GQA datasets [26]. Lastly, the LLaVA-Bench dataset scrutinizes LVLMs’ adaptability to challenging tasks and new domains through 24 diverse images and 60 questions, assessing robustness to various prompts [31]. Comprehensive details on these datasets are provided in the appendix 0.B.
Evaluation of LVLM Baselines: The effectiveness of the proposed method is evaluated across three contemporary LVLMs. Specifically, we apply our methodology to LLaVA-1.5 and InstructBLIP, both utilizing Vicuna 7B as their language decoder [9, 30]. Additionally, we assess Qwen-VL and Qwen-VL-Chat, constructed on the Qwen 7B backbone [3]. In addition to the LVLMs, we introduce two Supervised Fine-Tuning (SFT) baseline approaches. Firstly, LLaVA-SFT [43] serves as a benchmark for our comparison. This model is trained on a novel mixture of data, combining synthetic vision instruction tuning data [31] with existing high-quality human-annotated multi-modal data in a conversational format. Secondly, LLaVA-RLHF [43] is trained using a LLaVA-SFT-13b initialized reward model and leverages the FACT-RLHF algorithm on the mixed data. To ensure a robust comparison, we present averaged results calculated over three runs on all benchmarks. Due to unreported generation parameters of VCD [21] on some benchmarks, we report re-implemented scores on these benchmarks. For all contrastive decoding strategies, we utilize the same hyperparameters (, ). We execute all experiments on a machine equipped with an AMD EPYC 7542 32-Core Processor and Tesla-V100 (40G) instances. To ensure robustness and reliability, we repeat each experiment three times using different random seeds. It is noteworthy that, at the time of our experimentation, the previously employed GPT-4 API (gpt-4-0314), as utilized in established benchmarks [43, 30], has been deprecated. Consequently, we adopt the updated GPT-4 API (gpt-4-0613) and present the re-evaluated performance in this paper.
3.2 Systematic Evaluation
Impact of Post-hoc Debias and Debias Sampling Methods on LVLM Performance: We conduct a comprehensive analysis by combining various debiasing strategies, including both post-hoc debias methods and debias sampling strategies, to investigate their influence on LVLM truthfulness and reasoning abilities. The results obtained from the MME dataset are presented in Tab. 1. Our key observations can be summarized as follows: (1) Post-hoc debias methods significantly enhance model truthfulness by mitigating hallucination, with both None and Unk strategies exhibiting superior performance compared to the naive approach, particularly in mitigating object and attribute hallucination; (2) Debias sampling strategies strike a better balance, with VDD-None and VDD-Unk achieving competitive hallucination scores while consistently outperforming in reasoning scores; (3) Combining different post-hoc debiasing strategies (Both) does not yield significant benefits, whereas combining debias sampling strategies shows promising results; (4) Post-hoc debiasing methods operate independently from sampling methods such as VCD and our proposed debias sampling methods. For instance, VCD (u), which combines VCD sampling with Unk post-hoc debiasing, outperforms VCD alone. Notably, VCD exhibits subpar performance with the LLaVA-13B model, likely due to its sensitivity to image noise levels. In contrast, our model relies on text rather than images, resulting in both higher efficiency and robustness. Similar patterns can also be observed in other datasets such as POPE (see Fig. 6), where the proposed debiasing methods and sampling techniques consistently outperform the naive and other baseline methods. Additionally, we notice that employing image-based debiasing methods such as Noise/Zero does not consistently yield improvements. In our experiments, we observe a much larger variance and sensitivity to question type with such methods, further emphasizing the robustness of our image-free debiasing strategies.
| LLaVA-v1.5-7B | ||||||||
| Subsets | Naive | Post-Hoc Debias | VCD[21] | Debias Sampling | ||||
| None | Unk | Both | VDD-None | VDD-Unk | Both | |||
| MME-OH | 313.33 | 335.00 | 328.33 | 303.33 | 303.33 | 316.67 | 318.33 | 328.33 |
| MME-AH | 260.00 | 315.00 | 305.00 | 275.00 | 283.33 | 296.67 | 291.67 | 291.67 |
| MME-RE | 317.86 | 302.50 | 299.29 | 282.50 | 312.14 | 279.64 | 337.50 | 346.07 |
| Overall | 891.19 | 952.50 | 932.62 | 860.83 | 898.81 | 892.98 | 947.50 | 966.07 |
| LLaVA-v1.5-13B | ||||||||
| Subsets | Naive | VCD[21] | Unk | VDD-Unk | Mixed Strategy | |||
| VCD (u) | VCD (n) | VDD-None (u) | VDD-None (n) | |||||
| MME-OH | 291.67 | 316.67 | 343.33 | 351.67 | 338.33 | 333.33 | 348.33 | 348.33 |
| MME-AH | 280.00 | 253.33 | 320.00 | 308.33 | 318.33 | 318.33 | 325.00 | 320.00 |
| MME-RE | 357.14 | 284.29 | 277.86 | 361.43 | 279.29 | 281.79 | 267.50 | 269.29 |
| Overall | 928.81 | 854.29 | 941.19 | 1021.43 | 935.95 | 933.45 | 940.83 | 937.62 |
Effect of Decoding Strategies on LVLM Performance: In Tab. 2 (left), we conduct a comprehensive assessment of the LLaVA-v1.5 model’s performance across three distinct datasets—POPE-MSCOCO (average F1 across three subsets), MME (total score across 3 evaluation subsets), and MMMU (average ACC across 36 subsets)—by employing various decoding strategies and settings. The “Default” column denotes the baseline performance using default sampling parameters, while the “Sampling” columns exhibit results obtained through diverse decoding strategies, including temperature (Temp ), Top-, and Top-. The “Overall” column highlights the optimal performance achieved across all sampling strategies on all subsets. Notably, the application of the optimal sampling strategy leads to a substantial improvement in the overall model performance compared to default configurations. To underscore the generality of this phenomenon beyond LLaVA models, we extend our evaluation to include Qwen-VL and InstructBLIP in the POPE-MSCOCO benchmark. The results consistently demonstrate performance enhancements with refined sampling strategies across existing LVLMs. In Tab. 2 (right), we emphasize the superior performance of our proposed method, labeled as “Ours”, in comparison to alternative strategies. Our method surpasses fine-tuned models on specific instruction datasets, highlighting the effectiveness of our refined sampling strategies. Moreover, the 7B model, employing these refined sampling strategies, outperforms the 13B model with different fine-tuning strategies, including both supervised fine-tuning and RLHF. Additional detailed results for all subsets can be found in the appendix (Tabs. 6, 8, 9 and 10). Furthermore, a detailed comparison with additional baseline models on the POPE three subsets is provided in Tab. 11.
| Setting | Default | Sampling | Overall | |||
| Dataset | LLaVA-v1.5 | Temp | Top- | Top- | ||
| POPE MSCOCO | 7B | 79.5 | 84.0 | 84.0 | 84.0 | 84.1 |
| 13B | 77.2 | 84.2 | 84.2 | 84.1 | 84.2 | |
| Qwen-VL | 81.8 | 83.6 | 83.7 | 83.3 | 83.7 | |
| InstructBLIP | 78.9 | 79.8 | 84.5 | 79.9 | 84.5 | |
| MMMU | 7B | 34.4 | 43.4 | 40.3 | 40.8 | 44.7 |
| 13B | 36.3 | 43.2 | 43.4 | 43.1 | 45.8 | |
| MME-OH | 7B | 313.3 | 363.3 | 345.0 | 355.0 | 363.3 |
| 13B | 340.0 | 355.0 | 340.0 | 355.0 | 355.0 | |
| MME-AH | 7B | 260.0 | 323.3 | 301.7 | 325.0 | 330.0 |
| 13B | 303.3 | 321.7 | 311.7 | 321.7 | 321.7 | |
| MME-RE | 7B | 317.9 | 409.6 | 443.2 | 423.6 | 449.6 |
| 13B | 295.4 | 428.9 | 405.7 | 415.0 | 453.9 | |
| Adversarial | Accuracy | Precision | Recall | F1 |
| 7B[21] | 79.0 | 83.1 | 72.8 | 77.6 |
| 13B[43] | 67.2 | - | - | 74.7 |
| SFT-13B[43] | 82.3 | - | - | 81.1 |
| RLHF-13B[43] | 82.3 | - | - | 80.5 |
| 7B (Ours) | 83.5 | 89.3 | 76.2 | 82.2 |
| 13B (Ours) | 84.3 | 91.6 | 75.6 | 82.8 |
| Popular | Accuracy | Precision | Recall | F1 |
| 7B[21] | 81.9 | 88.9 | 72.8 | 80.1 |
| 13B[43] | 73.6 | - | - | 78.2 |
| SFT-13B[43] | 84.0 | - | - | 82.6 |
| RLHF-13B[43] | 83.9 | - | - | 81.8 |
| 7B (Ours) | 85.9 | 94.3 | 76.3 | 84.4 |
| 13B (Ours) | 86.1 | 95.3 | 75.9 | 84.5 |
Overall Performance: Upon reaching the final iteration of our proposed method, we initially identify the optimal sampling configuration for all backbones. Subsequently, we apply debias sampling methods. Simultaneously, for classification tasks, we implement our post-hoc debiasing method, utilizing both None and Unk inputs concurrently. As depicted in Fig. 7, it becomes evident that VCD fails to consistently improve all baseline models, particularly in reasoning tasks. In contrast, our proposed method demonstrates superior results across all aspects and backbone models. Finally, the proposed method is assessed on the LLava-Bench, and the results presented in Fig. 8 demonstrate that, for open-ended generation tasks, VCD does not surpass the default decoding configuration consistently, although it occasionally outperforms in the complex subset. In contrast, our proposed VDDconsistently exhibits improvements over the default decoding strategy. The right columns of the VDD section illustrate that, for generation tasks, model performance is consistently influenced by various decoding configurations. Unlike multi-choice QA tasks, which favor a low temperature (), we observe that generation tasks benefit from a higher or a larger value. Detailed results and results with larger backbones are shown in Tabs. 16 and 17. Qualitative examples in Sec. 0.C.1 demonstrate that VDD is more helpful and less hallucinated, generating more precise illustrations.
3.3 Other Findings and Analysis
Calibration is particularly advantageous when a model lacks confidence in its predictions. In Figure 9, we analyze various post-hoc debiasing methods. It is evident that the ’Naive’ approach performs poorly on instances where the model exhibits low confidence, which aligns with intuition. In contrast, our proposed post-hoc debiasing methods yield significant performance improvements for such cases. As prediction confidence increases, the model’s predictions become more reliable and accurate, resulting in diminishing performance gains. Even at high confidence levels, around 0.9-1.0, our proposed methods achieve comparable results without detrimental effects. Consequently, the overall performance of our proposed method surpasses that of the naive approach. Notably, both image-based debiasing methods demonstrate superior performance on instances with low confidence scores, indicating aggressive adjustments to prediction results. However, such aggressive adjustments may lead to poorer predictions for high-confidence samples. Therefore, we default to using image-free debiasing methods, as they consistently yield improvements.
Navigating Challenges in Debiasing LLM-Biased Benchmarks: As illustrated in Fig. 10, we unravel the intricate interplay between debiasing methods and benchmarks influenced by LLM biases, with Naive’s performance centered around textual information. The post-hoc debiasing methods proposed exhibit inconsistent enhancements on the MMMU benchmark. Upon meticulous examination, when we entirely remove vision information by replacing the input image with an empty string (None) or the <unk> identifier (Unk), the model maintains comparable or superior performance in specific subsets. In these scenarios, our post-hoc methods, crafted to alleviate biases in LVLMs, fall short of yielding optimal results. Especially in instances where pure LLMs outshine, the application of post-hoc methods may prove detrimental. It is crucial to emphasize that benchmarks tailored for LVLMs should prioritize reliance on input images rather than solely on textual content. Consequently, our methods can act as valuable indicators; when the proposed debiasing methods exhibit suboptimal performance, it signals that the benchmark may tilt more towards LLMs and may not be apt for evaluating LVLMs effectively. In such cases, the proposed sampling methods, VDD, also demonstrate inferiority compared to Naive, with an average accuracy of versus .
4 Concluding Remarks
We undertake a comprehensive exploration of biases and challenges associated with LVLMs, particularly focusing on their interaction with underlying LLMs. Our investigation exposes notable biases in LVLM-generated content, primarily influenced by the language priors ingrained in LLMs rather than the visual inputs. To address these biases, we introduce innovative debiasing strategies, including post-hoc debiasing methods and debias sampling techniques. Our experiments demonstrate the effectiveness of these strategies in mitigating hallucination and enhancing reasoning capabilities in LVLMs. Post-hoc debiasing methods, such as None and Unk, significantly improve model truthfulness, particularly when models lack confidence in their predictions. Moreover, debias sampling strategies, denoted as VDD, strike a balance by exhibiting competitive hallucination scores while consistently outperforming in reasoning tasks. The proposed strategies contribute to the reliability and applicability of LVLMs, addressing biases associated with language priors. Furthermore, our exploration into the impact of decoding configurations on LVLM performance reveals substantial improvements by refining sampling strategies. Optimal decoding configurations unleash the full potential of existing LVLMs, surpassing default configurations and raising concerns about the fairness of existing evaluations. As the field advances, addressing biases and refining evaluation methodologies are pivotal for harnessing the full potential of LVLMs in real-world applications. The Limitations and Future Work are shown in appendix 0.D.
Potential Negative Impact. Exclusively prioritizing debiasing efforts may inadvertently introduce a conservative stance, potentially stifling the creative and exploratory facets of LVLMs. Striking a delicate balance between mitigating biases and preserving the generative potential is imperative to avoid constraining the model’s versatility. Additionally, the observed instability of LVLMs in response to decoding configurations underscores the potential for manipulation. Adversarial adjustments to decoding configurations could exploit this vulnerability, leading the model to generate harmful or realignment contents.
Appendix 0.A Related Work
The landscape of Large Vision-Language Models (LVLMs) has undergone significant changes, initially rooted in BERT-based language decoders and later evolving to incorporate LLMs. LVLMs, heavily reliant on advanced LLMs like GPTs [34, 4], PaLM [7, 1], BLOOM [33], LLaMA [46, 47], Alpaca [44], Vicuna [6], and Mistral [18], exhibit enhanced capabilities and performance, particularly through end-to-end training techniques. Recent model developments, including Flamingo [2], PaLI [20], PaLM-E [10], BLIP-2 [24], InstructBLIP [32], Otter [22], MiniGPT-4 [56], mPLUG-Owl [49], LLaVA [23], and QWen-VL [3], bring unique perspectives to challenges such as scaling pre-training, enhancing instruction-following capabilities, and overcoming alignment issues between image and text modalities. However, it’s crucial to note that while high-capability LLMs empower LVLMs, they also introduce biases, contributing to prevalent hallucination problems and generating unreliable content.
Hallucination in LVLMs is redefined from the NLP community’s original concept of generating factual yet ungrounded content to encompass content not supported by the associated image [45, 17]. Several strategies have been explored to mitigate hallucinations in LVLMs. Initial efforts, geared towards small-scale VLMs, involved fine-grained modality alignment [38] and tackling statistical bias in object co-occurrence through data augmentation [19]. However, the distinctive behaviors of LVLMs pose a significant challenge, rendering existing methods impractical for generalization and scaling up [48]. Recent studies have addressed this challenge by introducing hallucination-targeted datasets for fine-tuning [13, 29], training a post-hoc revisor to generate less hallucinatory outputs [55], or adopting factually augmented Reinforcement Learning from Human Feedback (RLHF)[43]. While these interventions effectively address object hallucination in LVLMs, the associated human effort and computational costs underscore the need for a simpler yet efficient approach. A closely related method, VCD[21], manipulates the decoding distribution to reduce hallucination. In this paper, we demonstrate that VCD does not consistently bring improvement, and we propose a more efficient and effective method.
Addressing Biases in LVLMs. Recognizing the significant impact of language priors on Visual Question Answering (VQA) tasks [15, 57, 53], prior endeavors have primarily concentrated on bias mitigation within training sets. These initiatives entail the formulation of sophisticated training strategies and the construction of balanced datasets to tackle inherent biases. However, extending such methodologies LVLMs, which leverage pre-existing LLMs as text encoders, presents formidable challenges due to the intricate nature and resource-intensive process of retraining or fine-tuning. Moreover, the influence of language priors ingrained in utilized LLMs on LVLM performance remains uncertain. It is also noteworthy that certain works [8, 41] have delved into addressing biases in vision-language models. However, it is crucial to highlight that these works predominantly focus on social and group biases within vision-language models such as CLIP. Despite similarities in titles and overarching themes, their primary focus diverges substantially from our research objectives. Our investigation is uniquely centered on the nuanced biases inherent in LVLMs, providing a distinct perspective and contributions to the broader landscape of debiasing efforts in natural language understanding and computer vision.
Appendix 0.B Dataset Details
Perception, Knowledge, and Reasoning: The MMMU dataset [51] consists of 11.5K meticulously collected multimodal questions sourced from college exams, quizzes, and textbooks. It spans six core disciplines (Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, Tech & Engineering) and covers 30 subjects and 183 subfields. The dataset incorporates 30 highly diverse image types, including charts, diagrams, maps, tables, music sheets, and chemical structures. Unlike existing benchmarks, MMMU focuses on advanced perception and reasoning with domain-specific knowledge, challenging models to perform tasks similar to those encountered by experts.
Perception, Cognition, and Hallucination: The MME dataset [11] distinguishes itself by simultaneously appraising perception and cognition capabilities, encompassing tasks such as OCR, coarse-grained object recognition (encompassing existence, count, position, and color), and fine-grained object recognition (encompassing movie posters, celebrities, scenes, landmarks, and artworks). With a total of 14 subtasks, MME delivers a diverse and comprehensive evaluation, fulfilling the imperative for a thorough assessment of MLLMs across diverse modalities and cognitive domains. The existence and count subsets are applicable for object-level hallucination evaluation (MME-OH), while the position and color subsets are suitable for attribute-level hallucination (MME-AH) assessment [21]. Besides, we include four reasoning subsets: code reasoning, text translation, commonsense reasoning, and numerical calculation to evaluate the cognition ability (MME-RE).
Object Hallucination: POPE [26] introduces a streamlined method for evaluating object hallucination in Large Vision-Language Models (LVLMs). In this assessment framework, LVLMs are tasked with determining the presence of a specific object in a given image. The benchmark ensures a balanced ratio between queries probing existent and non-existent objects (i.e., 50% each). POPE incorporates three distinct sampling settings: random, popular, and adversarial, each creating negative samples in unique ways. In the random setting, objects absent from the image are chosen randomly, while the popular setting selects missing objects from a high-frequency pool. In the adversarial setting, priority is given to co-occurring objects not present in the image. The benchmark compiles data from three sources: MSCOCO [28], A-OKVQA [40], and GQA [16], utilizing 500 images from each dataset under each sampling setting. For each image, six questions are formulated, resulting in a total of 27,000 query-answer pairs from the development sets of these datasets. The evaluation focuses on four key metrics: Accuracy, Precision, Recall, and the F1 score.
Challenging Tasks and Generalizability to Novel Domains. LLaVA-Bench [31] is a dataset consisting of 24 images, spanning indoor and outdoor scenes, memes, paintings, and sketches, accompanied by a set of 60 questions designed to evaluate the capabilities of LVLMs. The dataset aims to assess the adaptability of LVLMs to challenging tasks and new domains. Each image is meticulously paired with a highly-detailed and manually-curated description, and questions are thoughtfully selected to encompass various contexts. Questions are categorized into conversation (simple QA), detailed description, and complex reasoning, aiming to evaluate the model’s robustness to diverse prompts.
Appendix 0.C Extend experimental results
1.
Person
2.
Bicycle
3.
Car
4.
Motorcycle
5.
Airplane
6.
Bus
7.
Train
8.
Truck
9.
Boat
10.
Traffic Light
11.
Fire Hydrant
12.
Stop Sign
13.
Parking Meter
14.
Bench
15.
Bird
16.
Cat
17.
Dog
18.
Horse
19.
Sheep
20.
Cow
21.
Elephant
22.
Bear
23.
Zebra
24.
Giraffe
25.
Backpack
26.
Umbrella
27.
Handbag
28.
Tie
29.
Suitcase
30.
Frisbee
31.
Skis
32.
Snowboard
33.
Sports Ball
34.
Kite
35.
Baseball Bat
36.
Baseball Glove
37.
Skateboard
38.
Surfboard
39.
Tennis Racket
40.
Bottle
41.
Wine Glass
42.
Cup
43.
Fork
44.
Knife
45.
Spoon
46.
Bowl
47.
Banana
48.
Apple
49.
Sandwich
50.
Orange
51.
Broccoli
52.
Carrot
53.
Hot Dog
54.
Pizza
55.
Donut
56.
Cake
57.
Chair
58.
Couch
59.
Potted Plant
60.
Bed
61.
Dining Table
62.
Toilet
63.
TV
64.
Laptop
65.
Mouse
66.
Remote
67.
Keyboard
68.
Cell Phone
69.
Microwave
70.
Oven
71.
Toaster
72.
Sink
73.
Refrigerator
74.
Book
75.
Clock
76.
Vase
77.
Scissors
78.
Teddy Bear
79.
Hair Drier
80.
Toothbrush
0.C.1 Qualitative examples
We present examples in Fig. 15 to illustrate that the default decoding strategy leads to significant hallucinations in some instances, resulting in a low GPT-4 evaluation score (2). In these cases, VDD proves highly effective in alleviating hallucinations and achieving higher scores. A similar pattern is observed in Fig. 16, indicating that merely scaling up the model cannot completely eliminate unrealistic patterns, while VDD consistently delivers superior performance. Examples in Fig. 17 further affirm the efficacy and accuracy of VDD, generating more precise illustrations.
Appendix 0.D Limitations and Future Work
While our study represents a significant step forward in addressing biases and enhancing the performance of Large Vision-Language Models (LVLMs), it is essential to recognize certain limitations and outline promising avenues for future research.
Diversifying Evaluation Scenarios
Limitation: Our current evaluation primarily revolves around existing open-sourced LVLMs and datasets, focusing predominantly on QA and classification tasks.
Future Work: Expanding our evaluation scope to encompass a broader range of applications and scenarios, such as medical diagnostics, accessibility, and domain-specific use cases. This will provide a more comprehensive understanding of LVLM performance in diverse real-world contexts.
Attention Mechanism Enhancement
Limitation: The attention mechanism in LVLMs sometimes exhibits harmful patterns, emphasizing special tokens rather than crucial vision or text tokens. Our current method does not directly manipulate the attention mechanism.
Future Work: Investigating techniques to enhance the attention mechanism, redirecting focus towards vision tokens and key text tokens. This exploration aims to refine the model’s interpretability and ensure more meaningful attention during inference.
Innovative Fusion Techniques
Limitation: Current fusion strategies for combining vision and language modalities may benefit from further innovation to enhance overall LVLM performance.
Future Work: Exploring novel techniques for multimodal fusion, including innovative attention mechanisms and feature fusion strategies. These advancements aim to improve the synergy between vision and language representations, ultimately boosting the model’s overall effectiveness.
In conclusion, addressing these limitations and venturing into promising future research directions will play a pivotal role in advancing the field of Large Vision-Language Models. Continued collaboration and exploration across these dimensions will contribute to the development of more robust, interpretable, and versatile LVLMs for a wide array of applications.
| Number | What is the number of cats in the image? |
| Relation | What is the relation between the cat and the sofa in the image? |
| Color | What is the color of the cat in the image? |
| Action | What is the cat doing in the image? |
| Shape | What is the shape of the cat in the image? |
| Method | Decoding Configuration | Github Link |
| VCD | Default | https://github.com/DAMO-NLP-SG/VCD |
| LLaVA-v1.5 | Greedy search | https://github.com/haotian-liu/LLaVA |
| InternLM | Beam (5) search | https://github.com/InternLM/InternLM-XComposer |
| PureMM | Greedy search | https://github.com/Q-MM/PureMM |
| INF-MLLM | Beam (5) search | https://github.com/infly-ai/INF-MLLM |
| LLaMA-Adapter | Top-p=0.75 | https://github.com/OpenGVLab/LLaMA-Adapter |
| Otter | Default | https://github.com/Luodian/Otter |
| Qwen-VL | Top-p=0.01 | https://github.com/QwenLM/Qwen-VL |
| LLaVA-v1.5-7B | InstructBLIP | ||||||||||
| Subset | Default | Temp | Top- | Top- | Overall | Subset | Default | Temp | Top- | Top- | Overall |
| Adversarial | 78.28 | 82.27 | 82.21 | 82.22 | 82.27 | Adversarial | 77.10 | 77.99 | 82.21 | 78.35 | 82.21 |
| Popular | 80.16 | 84.37 | 84.38 | 84.35 | 84.38 | Popular | 79.07 | 79.41 | 84.36 | 79.67 | 84.36 |
| Random | 80.00 | 85.49 | 85.51 | 85.51 | 85.51 | Random | 80.57 | 81.90 | 87.06 | 81.60 | 87.06 |
| Average | 79.48 | 84.05 | 84.03 | 84.03 | 84.05 | Average | 78.91 | 79.77 | 84.54 | 79.87 | 84.54 |
| LLaVA-v1.5-13B | Qwen-VL | ||||||||||
| Subset | Default | Temp | Top- | Top- | Overall | Subset | Default | Temp | Top- | Top- | Overall |
| Adversarial | 74.74 | 82.84 | 82.82 | 82.75 | 82.84 | Adversarial | 80.68 | 82.25 | 82.41 | 82.10 | 82.41 |
| Popular | 78.80 | 84.44 | 84.50 | 84.41 | 84.50 | Popular | 81.80 | 84.11 | 84.10 | 83.68 | 84.11 |
| Random | 78.20 | 85.17 | 85.25 | 85.10 | 85.25 | Random | 82.82 | 84.53 | 84.44 | 84.21 | 84.53 |
| Average | 77.25 | 84.15 | 84.19 | 84.09 | 84.20 | Average | 81.76 | 83.63 | 83.65 | 83.33 | 83.68 |
| LLaVA-v1.5-7B | |||||||||
| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||
| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||
| Object-level Hallucination | existence | 175.00 | 190.00 | 190.00 | 183.33 | 185.00 | 185.00 | 185.00 | 190.00 |
| count | 138.33 | 145.00 | 138.33 | 120.00 | 118.33 | 131.67 | 133.33 | 138.33 | |
| Attribute-level Hallucination | position | 128.33 | 145.00 | 140.00 | 130.00 | 128.33 | 131.67 | 133.33 | 126.67 |
| color | 131.67 | 170.00 | 165.00 | 145.00 | 155.00 | 165.00 | 158.33 | 165.00 | |
| Reasoning | commonsense | 102.86 | 95.00 | 94.29 | 105.00 | 102.14 | 92.14 | 120.00 | 113.57 |
| numerical | 60.00 | 70.00 | 67.50 | 45.00 | 92.50 | 70.00 | 77.50 | 70.00 | |
| translation | 92.50 | 50.00 | 50.00 | 47.50 | 70.00 | 72.50 | 92.50 | 75.00 | |
| code | 62.50 | 87.50 | 87.50 | 85.00 | 47.50 | 45.00 | 47.50 | 87.50 | |
| MME-OH | 313.33 | 335.00 | 328.33 | 303.33 | 303.33 | 316.67 | 318.33 | 328.33 | |
| MME-AH | 260.00 | 315.00 | 305.00 | 275.00 | 283.33 | 296.67 | 291.67 | 291.67 | |
| Reasoning | 317.86 | 302.50 | 299.29 | 282.50 | 312.14 | 279.64 | 337.50 | 346.07 | |
| LLaVA-v1.5-13B | |||||||||
| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||
| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||
| Object-level Hallucination | existence | 180.00 | 188.33 | 188.33 | 190.00 | 173.33 | 190.00 | 188.33 | 190.00 |
| 111.67 | 155.00 | 155.00 | 145.00 | 143.33 | 135.00 | 163.33 | 133.33 | ||
| Attribute-level Hallucination | position | 130.00 | 165.00 | 165.00 | 155.00 | 123.33 | 138.33 | 143.33 | 143.33 |
| 150.00 | 155.00 | 155.00 | 130.00 | 130.00 | 160.00 | 165.00 | 165.00 | ||
| Reasoning | commonsense | 112.14 | 115.00 | 112.86 | 122.14 | 134.29 | 115.71 | 116.43 | 111.43 |
| numerical | 70.00 | 32.50 | 32.50 | 42.50 | 25.00 | 57.50 | 57.50 | 37.50 | |
| translation | 100.00 | 42.50 | 62.50 | 52.50 | 67.50 | 70.00 | 102.50 | 77.50 | |
| code | 75.00 | 70.00 | 70.00 | 52.50 | 57.50 | 62.50 | 85.00 | 60.00 | |
| MME-OH | 291.67 | 343.33 | 343.33 | 335.00 | 316.67 | 325.00 | 351.67 | 323.33 | |
| MME-AH | 280.00 | 320.00 | 320.00 | 285.00 | 253.33 | 298.33 | 308.33 | 308.33 | |
| Reasoning | 357.14 | 260.00 | 277.86 | 269.64 | 284.29 | 305.71 | 361.43 | 286.43 | |
| LLaVA-v1.5-13B | Default | Temp | Top-k | Top-p | Overall | |
| Object-level Hallucination | existence | 185.0 | 195.0 | 185.0 | 195.0 | 195.0 |
| count | 155.0 | 160.0 | 155.0 | 160.0 | 160.0 | |
| Attribute-level Hallucination | position | 133.3 | 141.7 | 141.7 | 141.7 | 141.7 |
| color | 170.0 | 180.0 | 170.0 | 180.0 | 180.0 | |
| Reasoning | commonsense-reasoning | 127.9 | 136.4 | 125.7 | 135.0 | 136.4 |
| numerical-calculation | 42.5 | 77.5 | 72.5 | 70.0 | 77.5 | |
| text-translation | 77.5 | 125.0 | 92.5 | 105.0 | 125.0 | |
| code-reasoning | 47.5 | 90.0 | 115.0 | 105.0 | 115.0 | |
| All | 938.7 | 1105.6 | 1057.4 | 1091.7 | 1130.6 | |
| LLaVA-v1.5-7B | Default | Temp | Top-k | Top-p | Overall | |
| Object-level Hallucination | existence | 175.0 | 195.0 | 190.0 | 195.0 | 195.0 |
| count | 138.3 | 168.3 | 155.0 | 160.0 | 168.3 | |
| Attribute-level Hallucination | position | 128.3 | 143.3 | 131.7 | 150.0 | 150.0 |
| color | 131.7 | 180.0 | 170.0 | 175.0 | 180.0 | |
| Reasoning | commonsense-reasoning | 102.9 | 122.1 | 120.7 | 113.6 | 122.1 |
| numerical-calculation | 60.0 | 77.5 | 90.0 | 95.0 | 95.0 | |
| text-translation | 92.5 | 107.5 | 120.0 | 120.0 | 120.0 | |
| code-reasoning | 62.5 | 102.5 | 112.5 | 95.0 | 112.5 | |
| All | 891.2 | 1096.3 | 1089.9 | 1103.6 | 1143.0 | |
| Subsets | Default | Temp | Top- | Top- | Overall |
| Overall-Art and Design | 0.517 | 0.533 | 0.517 | 0.533 | 0.533 |
| Art | 0.533 | 0.600 | 0.567 | 0.567 | 0.600 |
| Art-Theory | 0.633 | 0.700 | 0.633 | 0.667 | 0.700 |
| Design | 0.533 | 0.600 | 0.633 | 0.600 | 0.633 |
| Music | 0.367 | 0.400 | 0.400 | 0.433 | 0.433 |
| Overall-Business | 0.240 | 0.293 | 0.267 | 0.260 | 0.293 |
| Accounting | 0.200 | 0.400 | 0.333 | 0.333 | 0.400 |
| Economics | 0.367 | 0.367 | 0.367 | 0.367 | 0.367 |
| Finance | 0.100 | 0.300 | 0.267 | 0.200 | 0.300 |
| Manage | 0.267 | 0.367 | 0.300 | 0.300 | 0.367 |
| Marketing | 0.267 | 0.300 | 0.233 | 0.300 | 0.300 |
| Overall-Science | 0.247 | 0.320 | 0.327 | 0.307 | 0.327 |
| Biology | 0.133 | 0.333 | 0.267 | 0.267 | 0.333 |
| Chemistry | 0.200 | 0.333 | 0.300 | 0.267 | 0.333 |
| Geography | 0.400 | 0.467 | 0.467 | 0.467 | 0.467 |
| Math | 0.267 | 0.467 | 0.367 | 0.400 | 0.467 |
| Physics | 0.233 | 0.367 | 0.333 | 0.367 | 0.367 |
| Overall-Health and Medicine | 0.333 | 0.360 | 0.367 | 0.347 | 0.367 |
| Basic-Medical-Science | 0.533 | 0.567 | 0.533 | 0.533 | 0.567 |
| Clinical-Medicine | 0.300 | 0.400 | 0.467 | 0.367 | 0.467 |
| Diagnostics-and-Laboratory-Medicine | 0.333 | 0.433 | 0.433 | 0.367 | 0.433 |
| Pharmacy | 0.367 | 0.500 | 0.367 | 0.400 | 0.500 |
| Public-Health | 0.133 | 0.333 | 0.400 | 0.333 | 0.400 |
| Overall-Humanities and Social Science | 0.492 | 0.533 | 0.500 | 0.508 | 0.533 |
| History | 0.433 | 0.533 | 0.433 | 0.433 | 0.533 |
| Literature | 0.767 | 0.800 | 0.767 | 0.800 | 0.800 |
| Sociology | 0.400 | 0.500 | 0.433 | 0.433 | 0.500 |
| Psychology | 0.367 | 0.467 | 0.467 | 0.500 | 0.500 |
| Overall-Tech and Engineering | 0.314 | 0.329 | 0.319 | 0.333 | 0.333 |
| Agriculture | 0.433 | 0.433 | 0.500 | 0.467 | 0.500 |
| Architecture-and-Engineering | 0.400 | 0.400 | 0.300 | 0.333 | 0.400 |
| Computer-Science | 0.200 | 0.300 | 0.300 | 0.367 | 0.367 |
| Electronics | 0.233 | 0.333 | 0.233 | 0.333 | 0.333 |
| Energy-and-Power | 0.400 | 0.500 | 0.400 | 0.400 | 0.500 |
| Materials | 0.267 | 0.433 | 0.367 | 0.400 | 0.433 |
| Mechanical-Engineering | 0.267 | 0.333 | 0.367 | 0.400 | 0.400 |
| Overall | 0.344 | 0.434 | 0.404 | 0.408 | 0.447 |
| Subsets | Default | Temp | Top- | Top- | Overall |
| Overall-Art and Design | 0.517 | 0.550 | 0.550 | 0.525 | 0.550 |
| Art | 0.600 | 0.633 | 0.633 | 0.600 | 0.633 |
| Art-Theory | 0.533 | 0.667 | 0.600 | 0.633 | 0.667 |
| Design | 0.633 | 0.667 | 0.700 | 0.667 | 0.700 |
| Music | 0.300 | 0.333 | 0.400 | 0.333 | 0.400 |
| Overall-Business | 0.227 | 0.280 | 0.307 | 0.293 | 0.307 |
| Accounting | 0.300 | 0.333 | 0.300 | 0.400 | 0.400 |
| Economics | 0.233 | 0.400 | 0.433 | 0.367 | 0.433 |
| Finance | 0.100 | 0.167 | 0.267 | 0.200 | 0.267 |
| Manage | 0.333 | 0.400 | 0.333 | 0.400 | 0.400 |
| Marketing | 0.167 | 0.300 | 0.333 | 0.300 | 0.333 |
| Overall-Science | 0.287 | 0.307 | 0.307 | 0.327 | 0.327 |
| Biology | 0.233 | 0.400 | 0.333 | 0.400 | 0.400 |
| Chemistry | 0.167 | 0.267 | 0.267 | 0.267 | 0.267 |
| Geography | 0.367 | 0.433 | 0.433 | 0.367 | 0.433 |
| Math | 0.367 | 0.433 | 0.433 | 0.433 | 0.433 |
| Physics | 0.300 | 0.300 | 0.367 | 0.500 | 0.500 |
| Overall-Health and Medicine | 0.393 | 0.420 | 0.413 | 0.407 | 0.420 |
| Basic-Medical-Science | 0.500 | 0.567 | 0.533 | 0.567 | 0.567 |
| Clinical-Medicine | 0.433 | 0.467 | 0.467 | 0.467 | 0.467 |
| Diagnostics-and-Laboratory-Medicine | 0.367 | 0.467 | 0.400 | 0.367 | 0.467 |
| Pharmacy | 0.267 | 0.333 | 0.333 | 0.267 | 0.333 |
| Public-Health | 0.400 | 0.467 | 0.533 | 0.433 | 0.533 |
| Overall-Humanities and Social Science | 0.533 | 0.567 | 0.567 | 0.575 | 0.575 |
| History | 0.567 | 0.600 | 0.633 | 0.600 | 0.633 |
| Literature | 0.700 | 0.767 | 0.767 | 0.733 | 0.767 |
| Sociology | 0.500 | 0.533 | 0.500 | 0.567 | 0.567 |
| Psychology | 0.367 | 0.400 | 0.433 | 0.467 | 0.467 |
| Overall-Tech and Engineering | 0.310 | 0.333 | 0.319 | 0.333 | 0.333 |
| Agriculture | 0.500 | 0.600 | 0.567 | 0.533 | 0.600 |
| Architecture-and-Engineering | 0.333 | 0.367 | 0.400 | 0.367 | 0.400 |
| Computer-Science | 0.200 | 0.267 | 0.333 | 0.333 | 0.333 |
| Electronics | 0.200 | 0.300 | 0.200 | 0.300 | 0.300 |
| Energy-and-Power | 0.400 | 0.500 | 0.500 | 0.467 | 0.500 |
| Materials | 0.367 | 0.433 | 0.433 | 0.467 | 0.467 |
| Mechanical-Engineering | 0.167 | 0.300 | 0.300 | 0.267 | 0.300 |
| Overall | 0.363 | 0.432 | 0.434 | 0.431 | 0.458 |
| Model | Accuracy | Precision | Recall | Yes | F1 |
| Adversarial | |||||
| LLaVA-v1.5-7B [21] | 78.96 | 83.06 | 72.75 | 77.57 | |
| LLaVA-v1.5-7B [43] | 62.70 | - | - | 83.00 | 72.00 |
| LLaVA-v1.5-13B [43] | 67.20 | - | - | 77.80 | 74.74 |
| LLaVA-SFT+7B [43] | 80.20 | - | - | 49.60 | 80.10 |
| LLaVA-SFT+13B [43] | 82.30 | - | - | 43.50 | 81.10 |
| LLaVA-RLHF-7B [43] | 80.70 | - | - | 44.00 | 79.50 |
| LLaVA-RLHF-13B [43] | 82.30 | - | - | 40.50 | 80.50 |
| LLaVA-v1.5-7B (Ours) | 83.52 | 89.25 | 76.22 | 42.70 | 82.22 |
| LLaVA-v1.5-13B (Ours) | 84.34 | 91.65 | 75.58 | 41.23 | 82.84 |
| Popular | |||||
| LLaVA-v1.5-7B [21] | 81.88 | 88.93 | 72.80 | - | 80.06 |
| LLaVA-v1.5-7B [43] | 68.40 | - | - | 77.90 | 75.30 |
| LLaVA-v1.5-13B [43] | 73.60 | - | - | 71.00 | 78.20 |
| LLaVA-SFT+7B [43] | 82.90 | - | - | 47.20 | 82.40 |
| LLaVA-SFT+13B [43] | 84.00 | - | - | 41.60 | 82.60 |
| LLaVA-RLHF-7B [43] | 83.30 | - | - | 41.80 | 81.80 |
| LLaVA-RLHF-13B [43] | 83.90 | - | - | 38.00 | 81.80 |
| LLaVA-v1.5-7B (Ours) | 85.87 | 94.32 | 76.33 | 40.47 | 84.38 |
| LLaVA-v1.5-13B (Ours) | 86.08 | 95.31 | 75.89 | 39.81 | 84.50 |
| Random | |||||
| LLaVA-v1.5-7B [21] | 83.29 | 92.13 | 72.80 | - | 81.33 |
| LLaVA-v1.5-7B [43] | 76.30 | - | - | 70.90 | 80.70 |
| LLaVA-v1.5-13B [43] | 73.70 | - | - | 72.30 | 78.80 |
| LLaVA-SFT+7B [43] | 86.10 | - | - | 45.50 | 85.50 |
| LLaVA-SFT+13B [43] | 82.30 | - | - | 43.50 | 81.10 |
| LLaVA-RLHF-7B [43] | 84.80 | - | - | 39.60 | 83.30 |
| LLaVA-RLHF-13B [43] | 85.20 | - | - | 38.40 | 83.50 |
| LLaVA-v1.5-7B (Ours) | 87.07 | 97.20 | 76.33 | 39.27 | 85.51 |
| LLaVA-v1.5-13B (Ours) | 86.88 | 97.29 | 75.87 | 38.99 | 85.25 |
| Dataset | Setting | Model | Decoding | Accuracy | Precision | Recall | F1 Score |
| MSCOCO | Random | LLaVA-7B | Naive | 83.29 | 92.13 | 72.80 | 81.33 |
| VCD | 87.73 | 91.42 | 83.28 | 87.16 | |||
| Ours | 90.03 | 97.36 | 79.13 | 88.79 | |||
| Qwen-VL | Naive | 84.73 | 95.61 | 72.81 | 82.67 | ||
| VCD | 88.63 | 94.64 | 81.91 | 87.81 | |||
| Ours | 89.20 | 92.55 | 85.27 | 88.80 | |||
| Popular | LLaVA-7B | Naive | 81.88 | 88.93 | 72.80 | 80.06 | |
| VCD | 85.38 | 86.92 | 83.28 | 85.06 | |||
| Ours | 88.57 | 94.32 | 79.33 | 87.00 | |||
| Qwen-VL | Naive | 84.13 | 94.31 | 72.64 | 82.06 | ||
| VCD | 87.12 | 91.49 | 81.85 | 86.40 | |||
| Ours | 87.67 | 90.94 | 83.67 | 87.15 | |||
| Adversarial | LLaVA-7B | Naive | 78.96 | 83.06 | 72.75 | 77.57 | |
| VCD | 80.88 | 79.45 | 83.29 | 81.33 | |||
| Ours | 84.00 | 90.03 | 76.47 | 82.70 | |||
| Qwen-VL | Naive | 82.26 | 89.97 | 72.61 | 80.37 | ||
| VCD | 84.26 | 85.84 | 82.05 | 83.90 | |||
| Ours | 85.23 | 89.06 | 80.33 | 84.61 | |||
| A-OKVQA | Random | LLaVA-7B | Naive | 83.45 | 87.24 | 78.36 | 82.56 |
| VCD | 86.15 | 85.18 | 87.53 | 86.34 | |||
| Ours | 88.13 | 93.27 | 82.20 | 88.24 | |||
| Qwen-VL | Naive | 86.67 | 93.16 | 79.16 | 85.59 | ||
| VCD | 89.22 | 90.77 | 87.32 | 89.01 | |||
| Ours | 89.43 | 89.15 | 89.80 | 89.47 | |||
| Popular | LLaVA-7B | Naive | 79.90 | 80.85 | 78.36 | 79.59 | |
| VCD | 81.85 | 78.60 | 87.53 | 82.82 | |||
| Ours | 86.57 | 90.07 | 82.20 | 85.95 | |||
| Qwen-VL | Naive | 85.56 | 90.44 | 79.53 | 84.63 | ||
| VCD | 87.85 | 88.10 | 87.53 | 87.81 | |||
| Ours | 88.47 | 87.02 | 88.07 | 88.54 | |||
| Adversarial | LLaVA-7B | Naive | 74.04 | 72.08 | 78.49 | 75.15 | |
| VCD | 74.97 | 70.01 | 87.36 | 77.73 | |||
| Ours | 80.23 | 79.09 | 82.20 | 80.61 | |||
| Qwen-VL | Naive | 79.57 | 79.77 | 79.23 | 79.50 | ||
| VCD | 81.27 | 77.79 | 87.53 | 82.38 | |||
| Ours | 82.33 | 82.21 | 82.53 | 82.37 | |||
| GQA | Random | LLaVA-7B | Naive | 83.73 | 87.16 | 79.12 | 82.95 |
| VCD | 86.65 | 84.85 | 89.24 | 86.99 | |||
| Ours | 88.97 | 93.72 | 83.53 | 88.33 | |||
| Qwen-VL | Naive | 80.97 | 88.07 | 71.64 | 79.01 | ||
| VCD | 85.59 | 86.88 | 83.84 | 85.33 | |||
| Ours | 87.37 | 85.50 | 90.00 | 87.69 | |||
| Popular | LLaVA-7B | Naive | 78.17 | 77.64 | 79.12 | 78.37 | |
| VCD | 80.73 | 76.26 | 89.24 | 82.24 | |||
| Ours | 85.13 | 87.37 | 82.07 | 84.68 | |||
| Qwen-VL | Naive | 75.99 | 78.62 | 71.40 | 74.84 | ||
| VCD | 81.83 | 80.45 | 84.09 | 82.23 | |||
| Ours | 82.43 | 77.85 | 91.00 | 83.90 | |||
| Adversarial | LLaVA-7B | Naive | 75.08 | 73.19 | 79.16 | 76.06 | |
| VCD | 76.09 | 70.83 | 88.75 | 78.78 | |||
| Ours | 81.81 | 81.70 | 82.07 | 81.77 | |||
| Qwen-VL | Naive | 75.46 | 77.92 | 71.07 | 74.33 | ||
| VCD | 80.01 | 77.86 | 83.85 | 80.75 | |||
| Ours | 78.80 | 73.35 | 90.00 | 81.21 |
| Dataset | Setting | Model | Decoding | Accuracy | Precision | Recall | F1 Score |
| MSCOCO | Random | LLaVA-v1.5-7B-SFT | Naive | 85.70 | 91.87 | 78.33 | 84.56 |
| VCD | 83.17 | 87.32 | 77.60 | 82.17 | |||
| Ours | 86.63 | 90.37 | 82.00 | 85.98 | |||
| LLaVA-v1.5-13B-SFT | Naive | 86.30 | 95.64 | 76.07 | 84.74 | ||
| VCD | 85.80 | 93.80 | 76.67 | 84.37 | |||
| Ours | 87.97 | 92.98 | 82.13 | 87.22 | |||
| Popular | LLaVA-v1.5-7B-SFT | Naive | 82.60 | 85.64 | 78.33 | 81.82 | |
| VCD | 79.73 | 81.06 | 77.60 | 79.29 | |||
| Ours | 83.23 | 84.79 | 79.93 | 82.82 | |||
| LLaVA-v1.5-13B-SFT | Naive | 83.97 | 90.34 | 76.07 | 82.59 | ||
| VCD | 83.9 | 89.63 | 76.67 | 82.64 | |||
| Ours | 84.93 | 88.64 | 80.13 | 84.50 | |||
| Adversarial | LLaVA-v1.5-7B-SFT | Naive | 79.6 | 80.33 | 78.40 | 79.35 | |
| VCD | 76.73 | 77.47 | 77.47 | 76.90 | |||
| Ours | 80.43 | 78.57 | 81.67 | 80.75 | |||
| LLaVA-v1.5-13B-SFT | Naive | 81.83 | 86.04 | 76.00 | 80.71 | ||
| VCD | 81.07 | 84.06 | 76.67 | 80.20 | |||
| Ours | 81.93 | 81.68 | 82.33 | 82.01 | |||
| A-OKVQA | Random | LLaVA-v1.5-7B-SFT | Naive | 84.33 | 85.03 | 83.33 | 84.18 |
| VCD | 82.10 | 81.00 | 83.87 | 82.41 | |||
| Ours | 84.40 | 84.54 | 84.20 | 84.37 | |||
| LLaVA-v1.5-13B-SFT | Naive | 87.2 | 90.73 | 82.87 | 86.62 | ||
| VCD | 85.90 | 88.06 | 83.07 | 85.49 | |||
| Ours | 87.37 | 88.95 | 85.33 | 87.10 | |||
| Popular | LLaVA-v1.5-7B-SFT | Naive | 79.53 | 77.45 | 83.33 | 80.28 | |
| VCD | 76.57 | 73.32 | 83.53 | 78.10 | |||
| Ours | 79.93 | 77.41 | 84.53 | 80.82 | |||
| LLaVA-v1.5-13B-SFT | Naive | 83.77 | 84.39 | 82.87 | 83.62 | ||
| VCD | 82.70 | 82.46 | 83.07 | 82.76 | |||
| Ours | 83.67 | 82.58 | 85.33 | 83.93 | |||
| Adversarial | LLaVA-v1.5-7B-SFT | Naive | 71.83 | 67.86 | 83.33 | 74.81 | |
| VCD | 70.53 | 66.28 | 83.60 | 73.94 | |||
| Ours | 72.60 | 68.25 | 84.53 | 75.52 | |||
| LLaVA-v1.5-13B-SFT | Naive | 76.97 | 74.12 | 82.87 | 78.25 | ||
| VCD | 76.17 | 72.83 | 83.47 | 77.79 | |||
| Ours | 76.80 | 72.89 | 85.33 | 78.62 | |||
| GQA | Random | LLaVA-v1.5-7B-SFT | Naive | 84.70 | 85.68 | 83.33 | 84.49 |
| VCD | 82.93 | 82.50 | 83.60 | 83.05 | |||
| Ours | 85.43 | 86.52 | 85.87 | 85.69 | |||
| LLaVA-v1.5-13B-SFT | Naive | 86.07 | 88.70 | 82.67 | 85.58 | ||
| VCD | 85.67 | 86.74 | 84.20 | 85.45 | |||
| Ours | 86.50 | 87.89 | 84.67 | 86.26 | |||
| Popular | LLaVA-v1.5-7B-SFT | Naive | 77.40 | 74.49 | 83.33 | 78.67 | |
| VCD | 74.57 | 70.81 | 83.60 | 76.67 | |||
| Ours | 78.53 | 75.44 | 84.87 | 79.87 | |||
| LLaVA-v1.5-13B-SFT | Naive | 81.10 | 80.16 | 82.67 | 81.39 | ||
| VCD | 80.20 | 77.96 | 84.20 | 80.96 | |||
| Ours | 82.03 | 80.09 | 85.67 | 82.60 | |||
| Adversarial | LLaVA-v1.5-7B-SFT | Naive | 73.27 | 69.24 | 83.73 | 75.80 | |
| VCD | 71.40 | 67.06 | 84.13 | 74.63 | |||
| Ours | 73.33 | 69.27 | 83.87 | 75.87 | |||
| LLaVA-v1.5-13B-SFT | Naive | 77.23 | 74.56 | 82.67 | 78.41 | ||
| VCD | 76.70 | 73.24 | 84.13 | 78.31 | |||
| Ours | 79.24 | 77.67 | 74.10 | 79.24 |
| Dataset | Setting | Model | Decoding | Accuracy | Precision | Recall | F1 Score |
| MSCOCO | Random | LLaVA-v1.5-13B | Naive | 86.80 | 97.42 | 75.60 | 85.14 |
| VCD | 86.93 | 95.48 | 77.53 | 85.58 | |||
| Ours | 89.20 | 95.72 | 82.07 | 88.37 | |||
| Qwen-VL-Chat | Naive | 86.93 | 98.34 | 75.13 | 85.19 | ||
| VCD | 87.23 | 96.89 | 76.93 | 85.77 | |||
| Ours | 89.23 | 96.16 | 81.73 | 88.36 | |||
| Popular | LLaVA-v1.5-13B | Naive | 86.10 | 95.70 | 75.60 | 84.47 | |
| VCD | 85.40 | 92.01 | 77.53 | 84.15 | |||
| Ours | 87.63 | 92.35 | 82.07 | 86.90 | |||
| Qwen-VL-Chat | Naive | 86.07 | 96.16 | 75.13 | 84.36 | ||
| VCD | 86.37 | 94.97 | 76.8 | 84.92 | |||
| Ours | 87.77 | 92.95 | 81.73 | 86.98 | |||
| Adversarial | LLaVA-v1.5-13B | Naive | 84.23 | 91.51 | 75.47 | 82.72 | |
| VCD | 83.23 | 87.45 | 77.60 | 82.23 | |||
| Ours | 84.97 | 87.07 | 82.13 | 84.53 | |||
| Qwen-VL-Chat | Naive | 84.17 | 91.77 | 75.07 | 82.58 | ||
| VCD | 83.73 | 89.41 | 76.53 | 82.47 | |||
| Ours | 84.50 | 86.47 | 81.80 | 84.07 | |||
| A-OKVQA | Random | LLaVA-v1.5-13B | Naive | 88.77 | 93.89 | 82.93 | 88.07 |
| VCD | 87.47 | 89.97 | 84.33 | 87.06 | |||
| Ours | 89.10 | 90.81 | 87.00 | 88.87 | |||
| Qwen-VL-Chat | Naive | 88.03 | 95.97 | 79.4 | 86.9 | ||
| VCD | 88.20 | 94.63 | 81.00 | 87.28 | |||
| Ours | 89.83 | 93.27 | 85.87 | 89.41 | |||
| Popular | LLaVA-v1.5-13B | Naive | 86.70 | 89.69 | 82.93 | 86.18 | |
| VCD | 85.33 | 86.05 | 84.33 | 85.19 | |||
| Ours | 88.00 | 88.78 | 87.00 | 87.88 | |||
| Qwen-VL-Chat | Naive | 87.7 | 95.20 | 79.40 | 86.59 | ||
| VCD | 87.63 | 93.19 | 81.20 | 86.78 | |||
| Ours | 89.40 | 92.34 | 85.93 | 89.02 | |||
| Adversarial | LLaVA-v1.5-13B | Naive | 80.17 | 78.58 | 82.93 | 80.70 | |
| VCD | 78.70 | 75.95 | 84.00 | 79.77 | |||
| Ours | 80.20 | 76.58 | 87.00 | 81.46 | |||
| Qwen-VL-Chat | Naive | 82.43 | 84.53 | 79.40 | 81.88 | ||
| VCD | 82.07 | 82.59 | 81.27 | 81.92 | |||
| Ours | 82.83 | 81.19 | 85.47 | 83.27 | |||
| GQA | Random | LLaVA-v1.5-13B | Naive | 89.53 | 93.73 | 84.73 | 89.01 |
| VCD | 88.87 | 91.11 | 86.13 | 88.55 | |||
| Ours | 89.80 | 91.46 | 87.80 | 89.59 | |||
| Qwen-VL-Chat | Naive | 87.60 | 93.32 | 81.00 | 86.72 | ||
| VCD | 87.03 | 91.42 | 81.73 | 86.31 | |||
| Ours | 88.67 | 88.67 | 88.67 | 88.67 | |||
| Popular | LLaVA-v1.5-13B | Naive | 86.40 | 87.66 | 84.73 | 86.17 | |
| VCD | 84.80 | 83.90 | 86.13 | 85.00 | |||
| Ours | 87.50 | 87.28 | 87.80 | 87.54 | |||
| Qwen-VL-Chat | Naive | 83.77 | 85.74 | 81.00 | 83.30 | ||
| VCD | 82.03 | 82.44 | 81.40 | 81.92 | |||
| Ours | 82.70 | 78.84 | 89.40 | 83.79 | |||
| Adversarial | LLaVA-v1.5-13B | Naive | 82.23 | 80.70 | 84.73 | 82.67 | |
| VCD | 80.97 | 78.07 | 86.13 | 81.90 | |||
| Ours | 82.47 | 79.34 | 87.80 | 83.35 | |||
| Qwen-VL-Chat | Naive | 82.53 | 83.56 | 81.00 | 82.26 | ||
| VCD | 81.53 | 81.58 | 81.47 | 81.52 | |||
| Ours | 81.23 | 77.19 | 88.67 | 82.53 |
| LLaVA-v1.5-7B | LLaVA-v1.5-13B | ||||||
| Split | Naive | VCD | Ours | Naive | VCD | Ours | |
| MME-OH | Existence | 175.0 | 185.0 | 190.0 | 180.0 | 173.3 | 195.0 |
| Count | 138.3 | 118.3 | 143.3 | 111.7 | 143.3 | 160.0 | |
| MME-AH | Position | 128.3 | 128.3 | 145.0 | 130.0 | 123.3 | 150.0 |
| Color | 131.7 | 155.0 | 165.0 | 150.0 | 130.0 | 180.0 | |
| MME-RE | Commonsense | 102.9 | 102.1 | 112.9 | 112.1 | 134.3 | 129.3 |
| Numerical | 60.0 | 92.5 | 57.5 | 70.0 | 25.0 | 90.0 | |
| Translation | 92.5 | 70.0 | 142.5 | 100.0 | 67.5 | 87.5 | |
| Code | 62.5 | 47.5 | 70.0 | 75.0 | 57.5 | 100.0 | |
| Overall | 891.2 | 898.8 | 1026.2 | 928.8 | 854.3 | 1091.8 | |
| LLaVA-v1.5-7B-SFT | LLaVA-v1.5-13B-SFT | ||||||
| Split | Naive | VCD | Ours | Naive | VCD | Ours | |
| MME-OH | Existence | 170.0 | 185.0 | 190.0 | 195.0 | 195.0 | 195.0 |
| Count | 120.0 | 140.0 | 151.7 | 165.0 | 170.0 | 170.0 | |
| MME-AH | Position | 66.7 | 60.0 | 96.7 | 86.7 | 81.7 | 101.7 |
| Color | 115.0 | 115.0 | 140.0 | 136.7 | 163.3 | 168.3 | |
| MME-RE | Commonsense | 121.4 | 120.0 | 128.6 | 125.7 | 119.3 | 132.9 |
| Numerical | 25.0 | 25.0 | 55.0 | 37.5 | 57.5 | 97.5 | |
| Translation | 50.0 | 50.0 | 77.5 | 67.5 | 80.0 | 107.5 | |
| Code | 52.5 | 65.0 | 75.0 | 60.0 | 52.5 | 95.0 | |
| Overall | 720.6 | 760.0 | 914.4 | 874.0 | 919.3 | 1067.9 | |
| Qwen-VL 7B | Qwen-VL-Chat | ||||||
| Split | Naive | VCD | Ours | Naive | VCD | Ours | |
| MME-OH | Existence | 155.0 | 165.0 | 165.0 | 173.3 | 165.0 | 185.0 |
| Count | 103.3 | 120.0 | 145.0 | 133.3 | 136.7 | 156.7 | |
| MME-AH | Position | 143.3 | 133.3 | 148.3 | 151.7 | 143.3 | 168.3 |
| Color | 175.0 | 185.0 | 190.0 | 170.0 | 170.0 | 185.0 | |
| MME-RE | Commonsense | 105.7 | 105.0 | 132.9 | 106.4 | 104.3 | 125.7 |
| Numerical | 60.0 | 57.5 | 77.5 | 45.0 | 37.5 | 62.5 | |
| Translation | 67.5 | 85.0 | 132.5 | 127.5 | 112.5 | 152.5 | |
| Code | 27.5 | 45.0 | 92.5 | 67.5 | 70.0 | 100.0 | |
| Overall | 837.4 | 895.8 | 1083.7 | 974.8 | 939.3 | 1135.7 | |
| LLava-v1.5-13B Absolute Score | |||||||||
| Subsets | Default | VCD[21] | VDD | Overall VDD | |||||
| Naive | =0.05 | =0.2 | =0.5 | 0.05 | 0.25 | ||||
| Complex | 75.70 | 76.50 | 75.30 | 77.30 | 75.80 | 76.20 | 76.00 | 77.00 | 77.30 |
| Conv | 67.70 | 66.20 | 70.20 | 69.80 | 68.30 | 68.80 | 70.00 | 69.80 | 70.20 |
| Detail | 60.00 | 56.80 | 61.00 | 62.50 | 62.70 | 62.70 | 63.30 | 62.00 | 63.30 |
| Full-Set | 67.80 | 66.50 | 68.80 | 69.90 | 68.90 | 69.20 | 69.80 | 69.60 | 70.27 |
| LLava-v1.5-13B Relative Score | |||||||||
| Subsets | Default | VCD[21] | VDD | Overall VDD | |||||
| Naive | =0.05 | =0.2 | =0.5 | 0.05 | 0.25 | ||||
| Complex | 89.00 | 89.70 | 88.60 | 89.60 | 88.90 | 89.30 | 90.50 | 89.50 | 90.50 |
| Conv | 78.40 | 77.50 | 81.30 | 80.60 | 79.10 | 79.10 | 80.20 | 80.90 | 81.30 |
| Detail | 74.70 | 70.20 | 75.60 | 77.20 | 78.30 | 79.00 | 79.20 | 77.80 | 79.20 |
| Full-Set | 80.80 | 79.30 | 81.90 | 82.50 | 82.20 | 82.50 | 83.30 | 82.90 | 83.67 |
| LLava-v1.5-7B Absolute Score | |||||||||
| Subsets | Default | VCD[21] | VDD | Overall VDD | |||||
| Naive | =0.05 | =0.2 | =0.5 | 0.05 | 0.25 | ||||
| Complex | 75.70 | 76.50 | 75.30 | 77.30 | 75.80 | 76.20 | 76.00 | 77.00 | 77.30 |
| Conv | 67.70 | 66.20 | 70.20 | 69.80 | 68.30 | 68.80 | 70.00 | 69.80 | 70.20 |
| Detail | 60.00 | 56.80 | 61.00 | 62.50 | 62.70 | 62.70 | 63.30 | 62.00 | 63.30 |
| Full-Set | 67.80 | 66.50 | 68.80 | 69.90 | 68.90 | 69.20 | 69.80 | 69.60 | 70.27 |
| LLava-v1.5-7B Relative Score | |||||||||
| Subsets | Default | VCD[21] | VDD | Overall VDD | |||||
| Naive | =0.05 | =0.2 | =0.5 | 0.05 | 0.25 | ||||
| Complex | 89 | 89.7 | 88.6 | 89.6 | 88.9 | 89.3 | 90.5 | 89.5 | 90.50 |
| Conv | 78.4 | 77.5 | 81.3 | 80.6 | 79.1 | 79.1 | 80.2 | 80.9 | 81.30 |
| Detail | 74.7 | 70.2 | 75.6 | 77.2 | 78.3 | 79 | 79.2 | 77.8 | 79.20 |
| Full-Set | 80.8 | 79.3 | 81.9 | 82.5 | 82.2 | 82.5 | 83.3 | 82.9 | 83.67 |

### v2

@@
-https://github.com/yfzhang114/LLaVA-Align Debiasing Large Visual Language Models
+https://github.com/yfzhang114/LLaVA-Align Debiasing Multimodal Large Language Models
 Abstract
@@
 Even in situations where the image is entirely noisy or absent, LVLMs confidently generate answers, indicating a bias towards the learned language patterns.
-This observation is illustrated in Fig.
-1, where querying an LVLM about the color of a non-existent fox produces confident yet unrealistic answers like brown, gray, or blue.
+This observation is illustrated in Fig.˜1, where querying an LVLM about the color of a non-existent fox produces confident yet unrealistic answers like brown, gray, or blue.
 To explore this bias further, we conduct a toy experiment using questions about color, shape, number, and relationships for 80 entities in the MSCOCO dataset.
 Our experimental baselines include Noise (fully noisy image and question input), None (no image provided), and Zeros/Ones (image with all zero/one values).
-Fig.
-1 displays the top 15 choices and corresponding probabilities of frequently occurring answers across these cases.
+Fig.˜1 displays the top 15 choices and corresponding probabilities of frequently occurring answers across these cases.
 Remarkably, even when no image is available or questions are about non-existent or meaningless images, current LVLMs tend to produce specific answers.
@@
 Our first approach introduces a “calibration” step to adjust the output distribution.
-Specifically, we assess the model’s inclination toward specific answers by inputting a dummy test with no image, as illustrated in Fig.
-1, where we replace the image with ’None’ or meaningless tensors.
+Specifically, we assess the model’s inclination toward specific answers by inputting a dummy test with no image, as illustrated in Fig.˜1, where we replace the image with ’None’ or meaningless tensors.
 We then determine the calibration parameters to ensure that the image-free input produces uniform scores for each answer without yielding overly confident outputs.
@@
 Our primary hypothesis is that existing evaluations are predominantly based on a default decoding setting, limiting the exploration of the model’s full capabilities.
-As extensively illustrated in Fig.
-2, diverse choices in generative configurations yield substantially different performance outcomes.
+As extensively illustrated in Fig.˜2, diverse choices in generative configurations yield substantially different performance outcomes.
 Notably, adjusting the generation temperature can lead to a notable increase in the average F1 score of POPE-MSCOCO, shifting from 76 to approximately 84.04.
@@
 Usually, an LLM is first pre-trained on a large, unlabeled text dataset, then an LVLM utilizes the LLM as the text encoder and incorporates vision representations in different ways [3, 23, 9, 12].
-Related work about LVLMs and Hallucination in LVLMs can be found in Appendix 0.A.
+Related work about LVLMs and Hallucination in LVLMs can be found in Appendix˜0.A.
 2.1 Exploring the Impact of LLM Biases on LVLMs
@@
 Furthermore, it remains unclear whether LVLM performance is affected by the language priors ingrained in their used LLMs.
-To assess the influence of LLM biases on LVLMs, we conduct toy experiments with five target questions detailed in Tab.
-4.
-For each entity of the MSCOCO dataset (appendix Tab.
-3), we design corresponding prompts and pair them with different types of vision inputs, including unrelated vision input such as fully noisy images (Noise) or entirely black or white images (Zero/One), or replace vision input with pure text, where we either remove all vision tokens (None) or replace them with a meaningless placeholder like </unk> (Unk).
-As illustrated in Fig.
-1 and in the appendix Fig.
-13, LVLMs exhibit biases towards specific answers, such as associating “white” with color-related questions, “round” with shape-related questions, and “many” with number-based questions.
+To assess the influence of LLM biases on LVLMs, we conduct toy experiments with five target questions detailed in Tab.˜4.
+For each entity of the MSCOCO dataset (appendix Tab.˜3), we design corresponding prompts and pair them with different types of vision inputs, including unrelated vision input such as fully noisy images (Noise) or entirely black or white images (Zero/One), or replace vision input with pure text, where we either remove all vision tokens (None) or replace them with a meaningless placeholder like </unk> (Unk).
+As illustrated in Fig.˜1 and in the appendix Fig.˜13, LVLMs exhibit biases towards specific answers, such as associating “white” with color-related questions, “round” with shape-related questions, and “many” with number-based questions.
 We further analyze attention mechanisms to explain why models heavily rely on LLM biases.
-Fig.
-4 and 14 demonstrate that LLM outputs tend to allocate more attention to text tokens.
+Fig.˜4 and 14 demonstrate that LLM outputs tend to allocate more attention to text tokens.
 Even for original image-text pairs where the number of image tokens exceeds that of text tokens, the sum of attention scores for text surpasses .
@@
 Most attention is directed towards special but uninformative tokens, rather than vision tokens or text sequences, such as the question itself.
-In open-ended generation tasks, we randomly select two questions from LLaVA-Bench (as shown in Fig.
-11) and examine the attention scores as more tokens were generated.
-As depicted in Fig.
-5, with the increasing length of generated text during open-ended tasks, there was a corresponding reduction in attention allocated to images, exacerbating the issue.
+In open-ended generation tasks, we randomly select two questions from LLaVA-Bench (as shown in Fig.˜11) and examine the attention scores as more tokens were generated.
+As depicted in Fig.˜5, with the increasing length of generated text during open-ended tasks, there was a corresponding reduction in attention allocated to images, exacerbating the issue.
 Consequently, the model becomes increasingly prone to generating content independently of input images, shedding light on the genesis of hallucinated content from an attention perspective.
@@
 2.2 Model Calibration
-Post-hoc Debiasing of Prediction Results (Fig.
-3 (left)).
+Post-hoc Debiasing of Prediction Results (Fig.˜3 (left)).
 In addressing the bias issue, we implement a “calibration” step for the model’s output probabilities using an affine transformation [36, 14, 52]: .
@@
 We then calculate the probability distribution .
-Visual Debias Decoding (VDD) (Fig.
-3 (Right)).
+Visual Debias Decoding (VDD) (Fig.˜3 (Right)).
 While post-hoc debias methods can be naturally extended to open-ended generation settings with the entire vocabulary as the label space, implementation faces challenges due to the large label space.
@@
 2.3 Impact of Decoding Configuration on Model Performance
-We observe that existing open-source LVLMs undergo evaluations using a variety of generation strategies, as outlined in Tab.
-5.
+We observe that existing open-source LVLMs undergo evaluations using a variety of generation strategies, as outlined in Tab.˜5.
 This variability in evaluation configuration introduces a potential source of performance variability, impacting the thorough exploration of LVLM capabilities.
@@
 Impact of Post-hoc Debias and Debias Sampling Methods on LVLM Performance: We conduct a comprehensive analysis by combining various debiasing strategies, including both post-hoc debias methods and debias sampling strategies, to investigate their influence on LVLM truthfulness and reasoning abilities.
-The results obtained from the MME dataset are presented in Tab.
-1.
+The results obtained from the MME dataset are presented in Tab.˜1.
 Our key observations can be summarized as follows: (1) Post-hoc debias methods significantly enhance model truthfulness by mitigating hallucination, with both None and Unk strategies exhibiting superior performance compared to the naive approach, particularly in mitigating object and attribute hallucination; (2) Debias sampling strategies strike a better balance, with VDD-None and VDD-Unk achieving competitive hallucination scores while consistently outperforming in reasoning scores; (3) Combining different post-hoc debiasing strategies (Both) does not yield significant benefits, whereas combining debias sampling strategies shows promising results; (4) Post-hoc debiasing methods operate independently from sampling methods such as VCD and our proposed debias sampling methods.
@@
 In contrast, our model relies on text rather than images, resulting in both higher efficiency and robustness.
-Similar patterns can also be observed in other datasets such as POPE (see Fig.
-6), where the proposed debiasing methods and sampling techniques consistently outperform the naive and other baseline methods.
+Similar patterns can also be observed in other datasets such as POPE (see Fig.˜6), where the proposed debiasing methods and sampling techniques consistently outperform the naive and other baseline methods.
 Additionally, we notice that employing image-based debiasing methods such as Noise/Zero does not consistently yield improvements.
@@
 | Overall | 928.81 | 854.29 | 941.19 | 1021.43 | 935.95 | 933.45 | 940.83 | 937.62 |
-Effect of Decoding Strategies on LVLM Performance: In Tab.
-2 (left), we conduct a comprehensive assessment of the LLaVA-v1.5 model’s performance across three distinct datasets—POPE-MSCOCO (average F1 across three subsets), MME (total score across 3 evaluation subsets), and MMMU (average ACC across 36 subsets)—by employing various decoding strategies and settings.
+Effect of Decoding Strategies on LVLM Performance: In Tab.˜2 (left), we conduct a comprehensive assessment of the LLaVA-v1.5 model’s performance across three distinct datasets—POPE-MSCOCO (average F1 across three subsets), MME (total score across 3 evaluation subsets), and MMMU (average ACC across 36 subsets)—by employing various decoding strategies and settings.
 The “Default” column denotes the baseline performance using default sampling parameters, while the “Sampling” columns exhibit results obtained through diverse decoding strategies, including temperature (Temp ), Top-, and Top-.
@@
 The results consistently demonstrate performance enhancements with refined sampling strategies across existing LVLMs.
-In Tab.
-2 (right), we emphasize the superior performance of our proposed method, labeled as “Ours”, in comparison to alternative strategies.
+In Tab.˜2 (right), we emphasize the superior performance of our proposed method, labeled as “Ours”, in comparison to alternative strategies.
 Our method surpasses fine-tuned models on specific instruction datasets, highlighting the effectiveness of our refined sampling strategies.
 Moreover, the 7B model, employing these refined sampling strategies, outperforms the 13B model with different fine-tuning strategies, including both supervised fine-tuning and RLHF.
-Additional detailed results for all subsets can be found in the appendix (Tabs.
-6, 8, 9 and 10).
-Furthermore, a detailed comparison with additional baseline models on the POPE three subsets is provided in Tab.
-11.
-| Setting | Default | Sampling | Overall | |||
-| Dataset | LLaVA-v1.5 | Temp | Top- | Top- | ||
-| POPE MSCOCO | 7B | 79.5 | 84.0 | 84.0 | 84.0 | 84.1 |
-| 13B | 77.2 | 84.2 | 84.2 | 84.1 | 84.2 | |
-| Qwen-VL | 81.8 | 83.6 | 83.7 | 83.3 | 83.7 | |
-| InstructBLIP | 78.9 | 79.8 | 84.5 | 79.9 | 84.5 | |
-| MMMU | 7B | 34.4 | 43.4 | 40.3 | 40.8 | 44.7 |
-| 13B | 36.3 | 43.2 | 43.4 | 43.1 | 45.8 | |
-| MME-OH | 7B | 313.3 | 363.3 | 345.0 | 355.0 | 363.3 |
-| 13B | 340.0 | 355.0 | 340.0 | 355.0 | 355.0 | |
-| MME-AH | 7B | 260.0 | 323.3 | 301.7 | 325.0 | 330.0 |
-| 13B | 303.3 | 321.7 | 311.7 | 321.7 | 321.7 | |
-| MME-RE | 7B | 317.9 | 409.6 | 443.2 | 423.6 | 449.6 |
-| 13B | 295.4 | 428.9 | 405.7 | 415.0 | 453.9 | |
+Additional detailed results for all subsets can be found in the appendix (Tabs.˜6, 8, 9 and 10).
+Furthermore, a detailed comparison with additional baseline models on the POPE three subsets is provided in Tab.˜11.
+| Setting | Default | Sampling | Overall | |||||
+| Dataset | LLaVA-v1.5 | Temp | Top- | Top- | ||||
+7B | 79.5 | 84.0 | 84.0 | 84.0 | 84.1 | ||
+| 13B | 77.2 | 84.2 | 84.2 | 84.1 | 84.2 | |||
+| Qwen-VL | 81.8 | 83.6 | 83.7 | 83.3 | 83.7 | |||
+| InstructBLIP | 78.9 | 79.8 | 84.5 | 79.9 | 84.5 | |||
+| MMMU | 7B | 34.4 | 43.4 | 40.3 | 40.8 | 44.7 | ||
+| 13B | 36.3 | 43.2 | 43.4 | 43.1 | 45.8 | |||
+| MME-OH | 7B | 313.3 | 363.3 | 345.0 | 355.0 | 363.3 | ||
+| 13B | 340.0 | 355.0 | 340.0 | 355.0 | 355.0 | |||
+| MME-AH | 7B | 260.0 | 323.3 | 301.7 | 325.0 | 330.0 | ||
+| 13B | 303.3 | 321.7 | 311.7 | 321.7 | 321.7 | |||
+| MME-RE | 7B | 317.9 | 409.6 | 443.2 | 423.6 | 449.6 | ||
+| 13B | 295.4 | 428.9 | 405.7 | 415.0 | 453.9 | |||
 | Adversarial | Accuracy | Precision | Recall | F1 |
@@
 Simultaneously, for classification tasks, we implement our post-hoc debiasing method, utilizing both None and Unk inputs concurrently.
-As depicted in Fig.
-7, it becomes evident that VCD fails to consistently improve all baseline models, particularly in reasoning tasks.
+As depicted in Fig.˜7, it becomes evident that VCD fails to consistently improve all baseline models, particularly in reasoning tasks.
 In contrast, our proposed method demonstrates superior results across all aspects and backbone models.
-Finally, the proposed method is assessed on the LLava-Bench, and the results presented in Fig.
-8 demonstrate that, for open-ended generation tasks, VCD does not surpass the default decoding configuration consistently, although it occasionally outperforms in the complex subset.
+Finally, the proposed method is assessed on the LLava-Bench, and the results presented in Fig.˜8 demonstrate that, for open-ended generation tasks, VCD does not surpass the default decoding configuration consistently, although it occasionally outperforms in the complex subset.
 In contrast, our proposed VDDconsistently exhibits improvements over the default decoding strategy.
@@
 Unlike multi-choice QA tasks, which favor a low temperature (), we observe that generation tasks benefit from a higher or a larger value.
-Detailed results and results with larger backbones are shown in Tabs.
-16 and 17.
-Qualitative examples in Sec.
-0.C.1 demonstrate that VDD is more helpful and less hallucinated, generating more precise illustrations.
+Detailed results and results with larger backbones are shown in Tabs.˜16 and 17.
+Qualitative examples in Sec.˜0.C.1 demonstrate that VDD is more helpful and less hallucinated, generating more precise illustrations.
 3.3 Other Findings and Analysis
@@
 Therefore, we default to using image-free debiasing methods, as they consistently yield improvements.
-Navigating Challenges in Debiasing LLM-Biased Benchmarks: As illustrated in Fig.
-10, we unravel the intricate interplay between debiasing methods and benchmarks influenced by LLM biases, with Naive’s performance centered around textual information.
+Navigating Challenges in Debiasing LLM-Biased Benchmarks: As illustrated in Fig.˜10, we unravel the intricate interplay between debiasing methods and benchmarks influenced by LLM biases, with Naive’s performance centered around textual information.
 The post-hoc debiasing methods proposed exhibit inconsistent enhancements on the MMMU benchmark.
-Upon meticulous examination, when we entirely remove vision information by replacing the input image with an empty string (None) or the <unk> identifier (Unk), the model maintains comparable or superior performance in specific subsets.
+Upon meticulous examination, when we entirely remove vision information by replacing the input image with an empty string (None∗) or the <unk> identifier (Unk∗), the model maintains comparable or superior performance in specific subsets.
 In these scenarios, our post-hoc methods, crafted to alleviate biases in LVLMs, fall short of yielding optimal results.
@@
 0.C.1 Qualitative examples
-We present examples in Fig.
-15 to illustrate that the default decoding strategy leads to significant hallucinations in some instances, resulting in a low GPT-4 evaluation score (2).
+We present examples in Fig.˜15 to illustrate that the default decoding strategy leads to significant hallucinations in some instances, resulting in a low GPT-4 evaluation score (2).
 In these cases, VDD proves highly effective in alleviating hallucinations and achieving higher scores.
-A similar pattern is observed in Fig.
-16, indicating that merely scaling up the model cannot completely eliminate unrealistic patterns, while VDD consistently delivers superior performance.
-Examples in Fig.
-17 further affirm the efficacy and accuracy of VDD, generating more precise illustrations.
+A similar pattern is observed in Fig.˜16, indicating that merely scaling up the model cannot completely eliminate unrealistic patterns, while VDD consistently delivers superior performance.
+Examples in Fig.˜17 further affirm the efficacy and accuracy of VDD, generating more precise illustrations.
 Appendix 0.D Limitations and Future Work
@@
 | Average | 77.25 | 84.15 | 84.19 | 84.09 | 84.20 | Average | 81.76 | 83.63 | 83.65 | 83.33 | 83.68 |
-| LLaVA-v1.5-7B | |||||||||
-| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||
-| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||
-| Object-level Hallucination | existence | 175.00 | 190.00 | 190.00 | 183.33 | 185.00 | 185.00 | 185.00 | 190.00 |
-| count | 138.33 | 145.00 | 138.33 | 120.00 | 118.33 | 131.67 | 133.33 | 138.33 | |
-| Attribute-level Hallucination | position | 128.33 | 145.00 | 140.00 | 130.00 | 128.33 | 131.67 | 133.33 | 126.67 |
-| color | 131.67 | 170.00 | 165.00 | 145.00 | 155.00 | 165.00 | 158.33 | 165.00 | |
-| Reasoning | commonsense | 102.86 | 95.00 | 94.29 | 105.00 | 102.14 | 92.14 | 120.00 | 113.57 |
-| numerical | 60.00 | 70.00 | 67.50 | 45.00 | 92.50 | 70.00 | 77.50 | 70.00 | |
-| translation | 92.50 | 50.00 | 50.00 | 47.50 | 70.00 | 72.50 | 92.50 | 75.00 | |
-| code | 62.50 | 87.50 | 87.50 | 85.00 | 47.50 | 45.00 | 47.50 | 87.50 | |
-| MME-OH | 313.33 | 335.00 | 328.33 | 303.33 | 303.33 | 316.67 | 318.33 | 328.33 | |
-| MME-AH | 260.00 | 315.00 | 305.00 | 275.00 | 283.33 | 296.67 | 291.67 | 291.67 | |
-| Reasoning | 317.86 | 302.50 | 299.29 | 282.50 | 312.14 | 279.64 | 337.50 | 346.07 | |
-| LLaVA-v1.5-13B | |||||||||
-| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||
-| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||
-| Object-level Hallucination | existence | 180.00 | 188.33 | 188.33 | 190.00 | 173.33 | 190.00 | 188.33 | 190.00 |
-| 111.67 | 155.00 | 155.00 | 145.00 | 143.33 | 135.00 | 163.33 | 133.33 | ||
-| Attribute-level Hallucination | position | 130.00 | 165.00 | 165.00 | 155.00 | 123.33 | 138.33 | 143.33 | 143.33 |
-| 150.00 | 155.00 | 155.00 | 130.00 | 130.00 | 160.00 | 165.00 | 165.00 | ||
-| Reasoning | commonsense | 112.14 | 115.00 | 112.86 | 122.14 | 134.29 | 115.71 | 116.43 | 111.43 |
-| numerical | 70.00 | 32.50 | 32.50 | 42.50 | 25.00 | 57.50 | 57.50 | 37.50 | |
-| translation | 100.00 | 42.50 | 62.50 | 52.50 | 67.50 | 70.00 | 102.50 | 77.50 | |
-| code | 75.00 | 70.00 | 70.00 | 52.50 | 57.50 | 62.50 | 85.00 | 60.00 | |
-| MME-OH | 291.67 | 343.33 | 343.33 | 335.00 | 316.67 | 325.00 | 351.67 | 323.33 | |
-| MME-AH | 280.00 | 320.00 | 320.00 | 285.00 | 253.33 | 298.33 | 308.33 | 308.33 | |
-| Reasoning | 357.14 | 260.00 | 277.86 | 269.64 | 284.29 | 305.71 | 361.43 | 286.43 | |
-| LLaVA-v1.5-13B | Default | Temp | Top-k | Top-p | Overall | |
-| Object-level Hallucination | existence | 185.0 | 195.0 | 185.0 | 195.0 | 195.0 |
-| count | 155.0 | 160.0 | 155.0 | 160.0 | 160.0 | |
-| Attribute-level Hallucination | position | 133.3 | 141.7 | 141.7 | 141.7 | 141.7 |
-| color | 170.0 | 180.0 | 170.0 | 180.0 | 180.0 | |
-| Reasoning | commonsense-reasoning | 127.9 | 136.4 | 125.7 | 135.0 | 136.4 |
-| numerical-calculation | 42.5 | 77.5 | 72.5 | 70.0 | 77.5 | |
-| text-translation | 77.5 | 125.0 | 92.5 | 105.0 | 125.0 | |
-| code-reasoning | 47.5 | 90.0 | 115.0 | 105.0 | 115.0 | |
-| All | 938.7 | 1105.6 | 1057.4 | 1091.7 | 1130.6 | |
-| LLaVA-v1.5-7B | Default | Temp | Top-k | Top-p | Overall | |
-| Object-level Hallucination | existence | 175.0 | 195.0 | 190.0 | 195.0 | 195.0 |
-| count | 138.3 | 168.3 | 155.0 | 160.0 | 168.3 | |
-| Attribute-level Hallucination | position | 128.3 | 143.3 | 131.7 | 150.0 | 150.0 |
-| color | 131.7 | 180.0 | 170.0 | 175.0 | 180.0 | |
-| Reasoning | commonsense-reasoning | 102.9 | 122.1 | 120.7 | 113.6 | 122.1 |
-| numerical-calculation | 60.0 | 77.5 | 90.0 | 95.0 | 95.0 | |
-| text-translation | 92.5 | 107.5 | 120.0 | 120.0 | 120.0 | |
-| code-reasoning | 62.5 | 102.5 | 112.5 | 95.0 | 112.5 | |
-| All | 891.2 | 1096.3 | 1089.9 | 1103.6 | 1143.0 | |
+| LLaVA-v1.5-7B | |||||||||||
+| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||||
+| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||||
+existence | 175.00 | 190.00 | 190.00 | 183.33 | 185.00 | 185.00 | 185.00 | 190.00 | ||
+| count | 138.33 | 145.00 | 138.33 | 120.00 | 118.33 | 131.67 | 133.33 | 138.33 | |||
+position | 128.33 | 145.00 | 140.00 | 130.00 | 128.33 | 131.67 | 133.33 | 126.67 | ||
+| color | 131.67 | 170.00 | 165.00 | 145.00 | 155.00 | 165.00 | 158.33 | 165.00 | |||
+| Reasoning | commonsense | 102.86 | 95.00 | 94.29 | 105.00 | 102.14 | 92.14 | 120.00 | 113.57 | ||
+| numerical | 60.00 | 70.00 | 67.50 | 45.00 | 92.50 | 70.00 | 77.50 | 70.00 | |||
+| translation | 92.50 | 50.00 | 50.00 | 47.50 | 70.00 | 72.50 | 92.50 | 75.00 | |||
+| code | 62.50 | 87.50 | 87.50 | 85.00 | 47.50 | 45.00 | 47.50 | 87.50 | |||
+| MME-OH | 313.33 | 335.00 | 328.33 | 303.33 | 303.33 | 316.67 | 318.33 | 328.33 | |||
+| MME-AH | 260.00 | 315.00 | 305.00 | 275.00 | 283.33 | 296.67 | 291.67 | 291.67 | |||
+| Reasoning | 317.86 | 302.50 | 299.29 | 282.50 | 312.14 | 279.64 | 337.50 | 346.07 | |||
+| LLaVA-v1.5-13B | |||||||||||
+| Subsets | Naive | Post-Hoc Debias | Debias Sampling | ||||||||
+| None | Unk | Both | VCD[21] | VDD-None | VDD-UNK | Both | |||||
+existence | 180.00 | 188.33 | 188.33 | 190.00 | 173.33 | 190.00 | 188.33 | 190.00 | ||
+| 111.67 | 155.00 | 155.00 | 145.00 | 143.33 | 135.00 | 163.33 | 133.33 | ||||
+position | 130.00 | 165.00 | 165.00 | 155.00 | 123.33 | 138.33 | 143.33 | 143.33 | ||
+| 150.00 | 155.00 | 155.00 | 130.00 | 130.00 | 160.00 | 165.00 | 165.00 | ||||
+| Reasoning | commonsense | 112.14 | 115.00 | 112.86 | 122.14 | 134.29 | 115.71 | 116.43 | 111.43 | ||
+| numerical | 70.00 | 32.50 | 32.50 | 42.50 | 25.00 | 57.50 | 57.50 | 37.50 | |||
+| translation | 100.00 | 42.50 | 62.50 | 52.50 | 67.50 | 70.00 | 102.50 | 77.50 | |||
+| code | 75.00 | 70.00 | 70.00 | 52.50 | 57.50 | 62.50 | 85.00 | 60.00 | |||
+| MME-OH | 291.67 | 343.33 | 343.33 | 335.00 | 316.67 | 325.00 | 351.67 | 323.33 | |||
+| MME-AH | 280.00 | 320.00 | 320.00 | 285.00 | 253.33 | 298.33 | 308.33 | 308.33 | |||
+| Reasoning | 357.14 | 260.00 | 277.86 | 269.64 | 284.29 | 305.71 | 361.43 | 286.43 | |||
+| LLaVA-v1.5-13B | Default | Temp | Top-k | Top-p | Overall | |||
+existence | 185.0 | 195.0 | 185.0 | 195.0 | 195.0 | ||
+| count | 155.0 | 160.0 | 155.0 | 160.0 | 160.0 | |||
+position | 133.3 | 141.7 | 141.7 | 141.7 | 141.7 | ||
+| color | 170.0 | 180.0 | 170.0 | 180.0 | 180.0 | |||
+| Reasoning | commonsense-reasoning | 127.9 | 136.4 | 125.7 | 135.0 | 136.4 | ||
+| numerical-calculation | 42.5 | 77.5 | 72.5 | 70.0 | 77.5 | |||
+| text-translation | 77.5 | 125.0 | 92.5 | 105.0 | 125.0 | |||
+| code-reasoning | 47.5 | 90.0 | 115.0 | 105.0 | 115.0 | |||
+| All | 938.7 | 1105.6 | 1057.4 | 1091.7 | 1130.6 | |||
+| LLaVA-v1.5-7B | Default | Temp | Top-k | Top-p | Overall | |||
+existence | 175.0 | 195.0 | 190.0 | 195.0 | 195.0 | ||
+| count | 138.3 | 168.3 | 155.0 | 160.0 | 168.3 | |||
+position | 128.3 | 143.3 | 131.7 | 150.0 | 150.0 | ||
+| color | 131.7 | 180.0 | 170.0 | 175.0 | 180.0 | |||
+| Reasoning | commonsense-reasoning | 102.9 | 122.1 | 120.7 | 113.6 | 122.1 | ||
+| numerical-calculation | 60.0 | 77.5 | 90.0 | 95.0 | 95.0 | |||
+| text-translation | 92.5 | 107.5 | 120.0 | 120.0 | 120.0 | |||
+| code-reasoning | 62.5 | 102.5 | 112.5 | 95.0 | 112.5 | |||
+| All | 891.2 | 1096.3 | 1089.9 | 1103.6 | 1143.0 | |||
 | Subsets | Default | Temp | Top- | Top- | Overall |